A Complete Bibliography of Publications in the Journal of the Society for Industrial and Applied Mathematics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

26 August 2024 Version 1.12

Title word cross-reference

∗ab^{x} + c [196]. ab^{x} + c [245]. a ∏ \frac{x−r_{i}}{x+s_{i}} [325]. Ax = λBx [222, 270]. Δu = u^{2} [142]. exp[−(sD)^{1/2}] [405]. \frac{∂^{2}u}{∂x^{2}} + \frac{∂^{2}u}{∂y^{2}} = \frac{∂u}{∂t} [27]. g [146]. H [492]. k [188]. kz^{2}K′_{n}(kz)K′_{n}(z) − n^{2}K_{n}(kz)K_{n}(z) [443]. L^{p} [387]. M [388]. N [394, 294, 187, 161, 322, 545, 96]. p^{m} [217]. T [550]. τ [39]. x′′ + cx′ + g(x) = Ef(t) [151].

-Algorithm [146, 550]. -Body [187]. -Frequency [545]. -Matrix [388]. -Method [39]. -Person [161]. -Theorem [492].

10 [397]. 11 [467, 555, 439, 440]. 13 [91, 558, 557].

288-293 [282, 270].

4 [556].

5 [373].

6 [132].

79 [91].

8 [238, 239].

9 [282, 225, 270, 296].

Absolute [485, 191]. Accelerating [237]. Acceleration [432, 94]. Accuracy [354]. Accurate [167]. Action [496]. Active [234, 322]. Adapted [111]. Additions [141]. Additive [119]. Adi [325]. Adjoint [82].

Aerofoils [382]. Affine [423]. Age [476]. Age- [476]. Aircraft [183]. Akad [91]. Algebraic [140]. Algorithm [205, 146, 91, 343, 116, 417, 442, 554, 550, 376]. Algorithms [327, 275, 175, 59, 245, 404]. Alias [158]. Alias-Free [158]. All-Integer [376]. Allocation [33, 286]. Almost [243]. Alternating [539, 154, 384, 469, 299, 543, 172, 278, 386]. Alternating-Direction-Implicit [384, 172, 278]. Alternatives [427]. Ambiguity [368]. Analog [354]. Analogue [147]. Analysis [337, 152, 70, 261, 14, 136, 543, 382, 553, 267]. Analytic [482]. Analytical [518, 187]. Anisotropic [358]. Annihilation [118]. Annulus [62]. Antena [396]. Antenna [307, 397]. Aperture [513]. Application [412, 117, 340, 228, 273, 77, 483, 75, 221, 459, 386, 325]. Applications [136, 150, 255, 405, 207, 47]. Applied [15, 482, 51, 76, 72]. Approach [66, 268, 253]. Approximate [104, 150, 120, 292, 305, 548]. Approximating [57]. Approximation [337, 482, 91, 65, 405, 196, 245, 266, 315, 253, 133, 109, 468, 325]. Approximations [194, 238, 387, 175, 370, 116, 415]. Arbitrary [260, 86, 360, 522]. Area [230]. Arguments [75]. Arising [20, 33, 421]. Arrival [351]. Aspects [52, 153, 69]. Assignment [395, 275, 59]. Assignments [409]. Associated [458]. Assumptions [65]. Asymptotic [254, 25, 45, 128, 556, 229, 416, 525, 167, 342, 523, 131]. Attraction [485]. Automata [499]. Automatic [1, 6, 12, 17, 21, 30, 37, 88, 282, 260, 141, 14]. Averaging [483]. Axial [531, 235]. Axial-Symmetric [531]. Axially [490, 4]. Axisymmetric [428].

Bairstow [349]. Ballistics [353, 392, 439]. Bar [226]. Barrier [76]. Base [307, 397]. Base-Driven [307, 397]. Based [162, 239, 267]. Basic [411]. be [381]. Beam [174, 429, 235]. Beams [449, 277, 341]. Behavior [254, 131]. Bending [240, 277, 298, 55, 436]. Bergman [61]. Bernoulli [160]. Bessel [373, 446, 435, 229, 58]. Best [188, 394, 116, 133]. Betting [527]. Between [44]. Bias [122]. Biased [76]. Bimatrix [430, 461]. Binary [503, 76, 552, 502]. Binomial [505]. Biology [10]. Biorthogonalization [80]. Biorthonormal [107]. Bipartite [327]. Block [54, 465]. Body [187]. Boltzmann [492]. Boolean [145, 350, 372, 423, 330, 331]. Bordered [293]. Bose [252]. Bose-Chaudhuri [252]. Boundary [289, 265, 489, 212, 321, 170, 42, 514, 271, 383, 459]. Bounds [412, 280, 334, 440, 137, 370, 416, 227, 68, 437, 109]. Branching [476, 374]. Buckling [515, 521]. Burst [552, 318]. Burst-Error [552].

Calculation [117, 210]. Calculus [249]. Can [381]. Capacitance [227]. Carlo [103]. Case [91]. Cauchy [166, 164]. cebysev [91]. Celestial [459]. Center [396]. Centered [184]. Certain [151, 281, 135, 179, 334, 440, 402, 256, 525, 403, 163, 42, 253, 69, 58, 373, 356, 36]. Chains [437, 294]. Channel [220]. Channels [224]. Characteristic [78, 157, 389, 222, 270]. Characteristics [111, 324]. Characterization [64, 451, 504]. Chaudhuri [252]. Chebyshev [387, 175, 196, 245, 266, 60, 325]. Checking [393]. Chemical [357]. Chinese [318]. Circle [302]. Circuit [335]. Circular [544, 513, 396, 304, 298, 375, 307, 397, 452]. Class [534, 265, 217, 518, 87, 170, 169, 324, 231, 519]. Classification [13, 423, 533]. Closed [67, 365]. Codes [217, 463, 252, 552, 163]. Coding [502]. Coefficient [489, 246, 361]. Coefficients [284, 39, 324, 438, 131].

Coherent [503]. Combinatorial [20, 71, 417, 306]. Combining [432]. Combustion [491]. Communication [288, 426]. Commutative [450, 474]. Comparison [408, 522, 60]. Competitive [243]. Complete [22, 366, 508, 558]. Complex [195, 101, 303]. Components [317]. Composite [427]. Computability [114]. Computation [74, 37, 88, 282, 260, 141, 79, 24, 457, 53, 102]. Computational [2]. Computations [277]. Computer [63, 24, 223]. Computers [152, 1, 6, 12, 17, 21, 30, 111]. Computing [85, 394, 38, 101, 14, 83]. Concentration [246, 361]. Concentration-Dependent [246]. Concepts [95]. Concerning [467, 91]. Condition [347, 380, 212, 441]. Conditions [547]. Conducting [513, 488, 410, 307, 397]. Conduction [354]. Conformal [47]. Conical [428]. Constant [310, 259]. Constants [24]. Constrained [38, 198]. Constraints [46, 315, 156, 241, 376]. Construction [350]. Containing [435, 522, 155]. Continuation [422]. Continuous [316, 464, 524, 519]. Contour [230]. Contraction [265]. Contributions [364]. Control [140, 14]. Convergence [320, 78, 44, 83, 7, 150, 186, 314, 116, 541, 178]. Convergent [355, 357]. Conversion [545]. Convex [205, 369, 197, 315, 352, 355, 356, 8]. Cooling [255]. Correcting [217]. Correction [88, 552, 318]. Corrections [141]. Corrector [328, 279, 383]. Correlation [263]. Correspondence [269]. Cosmical [310]. Cost [350, 264, 199]. Counter [76]. Counting [339, 458]. Crack [4]. Cracks [429]. Criteria [537, 211]. Criterion [234]. Critical [126]. Criticality [100, 132]. Cross [446, 301, 382]. Cross-Product [446]. Cross-Symmetry [301]. Crossings [508, 558]. Curtiss [441]. Curvature [232]. Curve [213, 77]. Curve-Fitting [213]. Cutting [197, 237]. Cutting-Plane [197]. Cyclic [463, 345, 83, 552]. Cylindrical [410, 452].

D. [555]. Damped [258]. Dams [283, 316]. Danilewski [101]. Data [140, 432, 63, 463, 189]. Data-Fitting [63]. Decision [479, 89, 40, 48]. Decision-Making [89]. Decoding [22, 106]. Decomposition [335, 257]. Defined [324, 535]. Deflating [206]. Deflation [200, 94]. Deflection [226]. Deformation [436]. Deformations [428]. Degeneracy [329]. Degree [362, 549, 510, 96]. Degree- [96]. Degrees [290, 323]. Delivery [551]. Delta [371, 155]. Demands [190]. Density [189]. Dependence [95]. Dependent [476, 246, 351]. Depending [128]. Depends [361]. Derivation [117, 506]. Derivatives [280]. Description [493, 219]. Design [335, 223]. Detailed [141]. Detection [192, 165, 177, 40, 48, 552]. Determinant [101]. Determination [322, 174, 189, 84]. Determine [223]. Determining [420]. Detonation [340]. Developing [82]. Diabatic [61]. Diagonal [72]. Difference [412, 347, 380, 233, 471, 212, 520, 518, 7, 460, 292, 147, 155, 271, 415, 109, 184]. Difference-Equation [292]. Differencing [210]. Differential [13, 144, 143, 309, 265, 516, 44, 328, 321, 501, 518, 402, 526, 170, 272, 421, 128, 299, 39, 460, 200, 26, 287, 403, 139, 342, 324, 438, 425, 130, 82, 129]. Differential-Difference [518]. Differentiation [207, 164, 359]. Diffraction [513]. Diffusion [246, 7, 421, 210, 361]. Digit [247]. Digital [122, 152, 1, 6, 12, 17, 21, 30, 111, 63, 24]. Dimension [445]. Dimensional [337, 340, 450]. Dimensions [271, 383]. Direction [539, 154, 384, 469, 299, 543, 172, 278, 386]. Dirichlet [142, 280, 419]. Discontinuous [401]. Discrete [70, 401, 150, 186, 532, 231, 178]. Discrete-Ordinate [186]. Discussion [333, 363]. Disk [396]. Disordered [97]. Distribution [281, 136, 410, 540, 547, 508, 558, 35, 523]. Distributions [89, 291, 247]. Doklady [91]. Domains [485]. Double [269, 277, 75]. Driven [307, 397]. Dual [203, 264, 454]. Duality [433, 352, 500, 519]. Due [36]. Duffing [363]. Durable [176]. Dynamic [544, 15, 46, 268, 190]. Dynamics [541].

Early [185]. Earth [385]. Echoes [165, 177]. Effect [160, 392, 436]. Effects [410]. Eigenfrequencies [444]. Eigenfunctions [57]. Eigenvalue [212, 99, 200, 530, 267]. Eigenvalues [117, 85, 282, 260, 101, 86, 83, 334, 440, 127, 420, 169, 457, 57, 84, 102, 82]. Eigenvectors [282, 260, 169, 102]. Elastic [449, 226, 104, 495]. Elastically [68]. Elasticity [555, 254, 390, 548]. Elastostatics [406]. Electrical [444]. Electromagnetic [513, 16]. Electronic [24]. Elements [467, 84]. Ellipsoidal [240]. Elliptic [524, 460, 26, 530, 382, 425, 436]. Emission [285, 214, 296]. Emulsion [168, 338]. Energy [496]. Engineering [10]. Enlarged [326]. Entries [377]. Entropy [542]. Equation [33, 144, 280, 419, 134, 389, 399, 398, 7, 256, 292, 498, 112, 287, 324, 276, 211]. Equations [215, 13, 388, 143, 91, 265, 516, 44, 233, 328, 471, 321, 454, 539, 520, 518, 402, 526, 127, 204, 170, 401, 506, 272, 193, 469, 431, 421, 450, 128, 120, 299, 107, 39, 460, 174, 26, 305, 403, 139, 342, 438, 47, 155, 425, 415, 130, 218, 184, 82, 129]. Equiconvergence [342]. Equilibrium [521, 430, 461, 171, 357]. Equipment [34, 195, 173, 176]. Equivalence [468]. Equivalent [400]. Erdélyi [556]. Errata [238, 132, 282, 439, 440, 270, 296, 397, 373, 557]. Erratum [225, 558, 239]. Error [148, 422, 217, 137, 370, 416, 552, 318, 109, 267]. Error-Correcting [217]. Error-Correction [552]. Errors [122, 14, 201]. Especially [56]. Essay [248]. Estimates [422, 477, 135, 522]. Estimating [432]. Estimation [300, 343, 522, 119]. Evaluation [11, 101, 326, 105, 166]. Evasion [434]. Even [221]. Every [294]. Exact [367]. Excitation [37, 88, 141]. Existence [353, 392, 439, 478]. Expansion [535, 480, 176]. Expansions [148, 73, 45, 170, 107, 556, 229, 416, 291]. Expectations [511]. Experience [2]. Experiment [313]. Explicit [121]. Explosion [531]. Exponential [272]. Exponential-like [272]. Exponentials [256, 266]. Expression [365]. Extended [386]. Extension [251, 185, 375]. Exterior [64]. External [302]. Extinction [476]. Extraction [40, 48]. Extrapolation [139].

Failure [317, 195]. Family [137]. Faulty [224]. Feedback [418, 286]. Field [396]. Fields [262, 163]. Films [24]. Filtering [192, 50, 3]. Filters [194, 238, 122, 231]. Finding [205, 179, 86, 431]. Finite [412, 283, 91, 262, 528, 212, 66, 171, 437, 163, 475, 147, 442, 554, 109]. Finite-Difference [212, 109]. Finite-State [437]. Fins [255, 191, 216, 259]. First [421, 370, 342, 271, 383]. First-Order [342, 271, 383]. Fit [320, 162, 239]. Fitting [63, 225, 208, 213, 77]. Fixed [246, 36]. Flat [297]. Fleet [183]. Flow [455, 507, 199, 61, 493, 220, 210, 297]. Flows [111, 244, 273]. Fluid [488]. Force [488]. Form [113, 309, 79, 484, 417]. Forms [136, 175]. Formula [405, 147, 221]. Formulae [337, 82]. Formulas [350, 279, 336, 480, 330, 331]. Formulation [140, 89]. Foundation [226]. Four [484]. Four-Person [484]. Fourier [556, 25, 45]. Fraction [405]. Fractional [490, 374]. Fracture [429]. Fredholm [47]. Free [304, 158, 496]. Frequencies [174, 223, 68]. Frequency [517, 545]. Fuel [236]. Function [362, 179, 448, 435, 484, 368, 404, 496].

Functional [33, 281, 401, 335, 139]. Functionals [464]. Functions [145, 148, 446, 399, 225, 208, 372, 423, 229, 35, 475, 324, 535, 221, 155, 359, 58, 373, 356]. Fundamental [517, 453, 401].

Game [456, 348]. Games [161, 243, 430, 484, 461, 171, 8]. Gas [493]. Gaussian [427, 119]. Gegenbauer [378]. General [50, 22, 79, 423, 450, 417, 147, 313, 382, 276, 231]. Generalization [234, 349, 61, 535, 480, 250, 95]. Generalized [364, 447, 173, 273, 213, 529, 305, 546]. Generation [63, 407]. Geometric [314, 531]. Geometry [462, 309, 104]. Gibbs [496]. Goursat [180]. Gradient [161, 198, 156, 241, 553]. Gradients [259]. Graph [290, 323]. Graphs [467, 455, 507, 262, 327, 269, 508, 558, 106, 499]. Gravitational [311]. Gravity [392]. Greatest [74]. Gregory [115]. Grids [289]. Griffin [555]. Groups [423, 89]. Growth [287].

Hakimi [467]. Hankel [554]. Harmonic [308, 441]. Having [84]. Heat [354, 191]. Heated [159]. Hermite [73, 147, 473]. Hermitian [85, 83]. High [111]. High-Speed [111]. Highly [488]. Hilbert [462, 391, 305]. Hill [112]. History [32]. Hölder [451]. Hole [398, 375]. Hydrodynamic [210]. Hyperbolic [265, 44, 471, 299, 271, 383, 109, 184]. Hyperbolic-Parabolic [265]. Hypergeometric [448]. Hypothesis [427].

Identities [117, 505]. II [181, 6, 323, 392, 19, 150, 48, 241, 331]. III [12, 23]. Impervious [424]. Implicant [248]. Implicit [27, 233, 384, 543, 210, 172, 278, 386]. Implicitly [535]. Imprimitive [389]. In-Line [223]. Incidence [529]. Including [410, 330, 331]. Incompatible [91]. Incomplete [233]. Independence [96]. Independent [225, 208, 294, 359]. Industrial [1, 6, 12, 17, 21, 30]. Inequalities [517, 526, 401]. Infinite [226, 159, 435, 375, 361]. Infinitely [513]. Influence [89]. Inhomogeneous [415]. Initial [347, 380, 235]. Input [545]. Integer [376]. Integers [290, 323]. Integral [399, 454, 378, 526, 127, 61, 107, 47]. Integrals [11, 25, 45, 435, 556, 166, 164]. Integrating [92, 181]. Integration [134, 27, 490, 110, 137, 201, 279, 62, 139, 480, 359, 182]. Integro [144]. Integro-Differential [144]. Integrodifferential [498]. Inter [351]. Inter-Arrival [351]. Internal [302]. Internal-External [302]. Interpolation [308, 75, 147, 221, 441, 149, 473]. Interpretation [480]. Intervals [112]. Introduction [31, 9, 19, 23, 16]. Introductory [124]. Invariant [231]. Inventory [190, 413]. Inverse [447, 529, 69, 504]. Inverses [364, 546]. Inversion [455, 228, 80, 378, 105, 442, 554, 118]. Inverting [67, 87]. Involving [45, 321, 399, 256, 556, 130, 58, 373]. Irreducible [510]. Isoenergetic [111]. Iteration [54, 233, 284, 72, 154, 172, 278, 386, 553, 325]. Iterations [28, 193]. Iterative [142, 215, 511, 81, 187, 359, 60]. IV [17, 31].

J [238, 132, 467, 555, 282, 225, 439, 440, 556, 558, 270, 296, 397, 373, 239, 557]. J. [256]. Jacobi [282, 85, 260, 345, 83]. Jacobi-like [282, 260]. Jeep [551]. Jitter [346].

Karman [487]. Katz [338]. Kernels [57]. Kilter [199]. Kinds [317]. Kinetic [493]. Kolmogorov [228]. Kramer [408]. Kron [333].

L [467]. Laboratory [24]. Lag [413]. Lagrange [462, 38, 535, 480, 473]. Lagrange-Hermite [473]. Laminar [297]. Lanczos [86]. Langenhop [512]. Laplace [419, 398, 547, 504]. Large [215, 396, 81, 129]. Large-Scale [215, 81]. Largest [162, 239, 523]. Lateral [235]. Lattice [97, 499]. Lattice-Ordered [499]. Lattices [56]. Laurent [131]. Law [195]. Laws [496]. Layer [489]. Least [122, 343]. Least-Squares [343]. Legendre [399, 407]. Length [230]. Like [471, 220, 282, 260, 272]. Limitations [232]. Limited [5, 424]. Line [223]. Linear [215, 50, 234, 322, 140, 70, 419, 143, 91, 262, 44, 273, 358, 290, 323, 423, 72, 400, 2, 28, 272, 77, 193, 61, 81, 128, 198, 175, 39, 129, 509, 552, 417, 156, 411, 324, 522, 438, 541, 500, 95, 415, 235, 130, 407, 519, 319, 536, 557, 82, 329, 8, 108]. Linear-Convex [8]. Linearly [361]. Lines [488]. Liouville [412, 284, 218]. Location [444, 470]. Locus [251, 365]. Logarithmic [45, 556]. Loopless [394]. Lower [227, 68]. Lyapunov [537, 250].

Machine [114]. Magnetic [488]. Making [89]. Mapping [47]. Marginal [319]. Markov [144, 294, 458]. Mass [496, 191, 216, 259]. Masses [223]. Math [91]. Mathematical [52, 295, 153, 433, 333, 89, 486, 494]. Mathematics [482]. Mathieu [276]. Matric [157]. Matrices [100, 132, 293, 85, 455, 477, 101, 86, 99, 83, 334, 440, 80, 67, 87, 206, 169, 546, 457, 84, 442, 554, 504, 474]. Matrix [117, 74, 254, 388, 447, 516, 362, 389, 260, 228, 79, 66, 529, 510, 547, 174, 105, 94, 249, 219, 102, 118]. Maximally [182]. Maximization [5]. Maximum [295, 551]. Mean [406, 148, 360]. Means [35]. Measure [475]. Measurements [332]. Mechanical [223]. Mechanics [459]. Mechanism [521]. Medium [159, 531, 361]. Membrane [517, 55]. Membranes [68].

Metal [24]. Method [54, 103, 511, 74, 340, 432, 78, 471, 282, 260, 111, 465, 25, 539, 199, 98, 86, 333, 420, 150, 186, 197, 384, 469, 431, 251, 365, 61, 366, 198, 39, 303, 483, 94, 305, 472, 139, 222, 270, 42, 167, 156, 241, 210, 306, 121, 60, 218, 178, 82, 237]. Methods [192, 412, 85, 27, 328, 101, 240, 110, 345, 83, 67, 87, 279, 65, 16, 299, 543, 292, 60]. Minima [38]. Minimal [350, 199, 497, 330, 331]. Minimal-Cost [199]. Minimality [22]. Minimization [253]. Minimize [527, 201]. Minimizing [404, 356]. Minimum [264, 65, 411, 236, 191, 216, 259]. Minimum-Fuel [236]. Minimum-Mass [191, 259]. Minimum-Time [236]. Mixed [432, 417, 133, 271, 383]. Mixing [489, 491]. Mode [277]. Model [168, 413, 338, 278, 386]. Models [10, 312]. Modes [223, 277]. Modifications [507]. Modified [229]. Modulation [545]. Modules [503]. Monotonicity [179]. Monte [103]. Motion [506, 232]. Multi [20, 244, 463, 72, 279, 209]. Multi-Diagonal [72]. Multi-Phase [463]. Multi-Stage [20]. Multi-Step [279]. Multi-Terminal [244]. Multi-Weighted [209]. Multidimensional [56]. Multiple [445, 505, 230, 29, 339, 377, 318]. Multiple-Burst [318]. Multiple-Valued [230]. Multipliers [462, 38]. Multipole [311]. Multistage [418, 348, 286]. Multivariate [75].

N.S [91]. Natural [274, 174, 223]. Nauk [91]. Navier [450]. Near [396, 385]. Nearly [221]. Nearly-Odd [221]. Necessary [347, 380, 547]. Neighboring [126]. Nerve [37, 88, 141]. Nets [22, 106]. Network [394, 244, 273, 426]. Networks [444, 288, 242, 371, 219]. Neumann [340]. Neutrices [125]. Neutron [300]. Newton [270, 117, 222]. no [238, 132, 467, 555, 282, 225, 439, 440, 556, 558, 270, 296, 397, 373, 239, 557]. Noise [73, 165, 177, 107, 40, 48, 424, 158, 119]. Non [85, 78, 526]. Non-Convergence [78]. Non-Hermitian [85]. Non-Oscillation [526]. Nonhomogeneous [517, 531]. Nonlinear [515, 328, 281, 99, 353, 392, 439, 478, 518, 431, 343, 460, 498, 428, 403, 156, 241, 90, 237]. Nonnegative [100, 132]. Nonsingularity [441]. Nonsymmetric [520]. Nonterminating [28]. Nonuniform [289]. Nonunique [521]. Normal [136, 223]. Norms [451]. Note [467, 92, 181, 388, 91, 453, 369, 344, 288, 225, 208, 464, 368, 556, 229, 55, 75, 96]. Notes [41]. Number [51, 372, 381]. Numbers [409, 502]. Numerical [300, 92, 181, 289, 555, 419, 143, 265, 134, 27, 44, 202, 240, 390, 110, 431, 256, 26, 62, 139, 210, 189, 341, 166, 359, 182, 297]. Nyquist [234].

Oblate [486]. Observational [312]. Obstacle [220]. Odd [221]. One [445, 340]. One-Dimensional [340]. Only [191]. Operation [29]. Operator [254, 464, 469, 61, 305]. Operators [92, 181, 524, 200]. Optical [24]. Optimal [283, 316, 479, 395, 138, 275, 409, 190, 286, 502]. Optimization [123, 418, 414]. Optimizing [511, 532]. Optimum [393, 317, 194, 238, 434, 65, 533, 545, 278, 224]. Optional [413]. Orbit [367, 483]. Orbits [486, 385]. Order [347, 402, 7, 421, 200, 287, 342, 294, 271, 383, 291, 129]. Ordered [502, 499]. Ordering [72, 204]. Ordinary [321, 272, 421, 128, 200, 139, 82]. Ordinate [150, 186, 178]. Ornstein [281]. Orr [134]. Orthogonal [63, 381, 267]. Oscillation [526]. Oscillations [525]. Oscillator [258]. Oscillators [478]. Osculatory [147]. Out-of-Kilter [199]. Overrelaxation [465, 305, 60]. p [91]. Pair [342]. Paper [555, 91, 256, 115]. Parabolic [265, 44, 233, 328, 539, 120, 26, 376]. Parallel [467]. Parameter [128, 119, 129]. Parameters [300, 322, 151, 343, 522, 278]. Parametric [170]. Part [392, 104, 156, 241, 382]. Partial [13, 265, 170, 460, 425]. Partition [484, 475]. Partitioned [447, 546]. Past [528]. Paths [394, 538]. Pattern [307, 397, 481]. Peg [398]. Perceptron [314]. Percolation [379, 374]. Periodic [151, 478, 501, 56, 324, 438]. Periodicity [534]. Persistence [160]. Person [161, 484]. Perturbation [226, 541]. Perturbational [276]. Perturbations [321]. Phase [25, 126, 463, 218]. Photographs [326]. Pitfalls [53]. Placed [381]. Planar [467, 159, 367, 483, 62]. Plane [513, 555, 242, 79, 390, 104, 197, 232, 237]. Planet [486]. Plate [304, 375, 297]. Plates [515, 68, 298]. Point [64, 137, 542, 40, 48, 370, 416, 531, 514, 130, 459]. Points [64, 126, 430, 461, 171]. Pointwise [186]. Poisson [280]. Polar [385]. Policies [283, 188, 479]. Policy [34, 190]. Polya [71]. Polyhedra [538]. Polyhedral [205]. Polymerization [168, 338]. Polynomial [157, 549, 303, 163, 441, 84, 291]. Polynomials [308, 135, 63, 60]. Position [476, 432]. Position-Dependent [476]. Positive [74, 547, 504]. Positive-Real [547, 504]. Potential [47]. Potentials [490, 36]. Power [94]. pp [91]. Practical [11]. Prediction [192, 50, 528]. Predictor [328, 383]. Predictor-Corrector [328, 383]. Predictors [194, 238]. Presence [36]. Present [114]. Prime [248]. Principle [509]. Priority [533]. Probabilistic [371]. Probabilities [476]. Probability [527, 10, 285, 437]. Problem [142, 412, 15, 46, 467, 280, 419, 91, 153, 395, 212, 98, 275, 284, 302, 180, 255, 97, 32, 367, 360, 533, 551, 222, 270, 187, 530, 69, 363, 271, 383, 121, 459, 278, 386, 357, 481]. Problems [192, 20, 289, 347, 380, 555, 73, 422, 64, 321, 199, 390, 99, 268, 170, 104, 150, 421, 354, 200, 483, 59, 18, 370, 416, 4, 42, 253, 47, 514, 519, 218, 257, 548]. Procedure [179, 213, 359, 355, 357]. Procedures [393, 137, 201]. Process [161, 144, 479, 281, 349, 284, 190, 542, 458]. Processes [43, 49, 20, 476, 418, 491, 261, 379, 374, 427, 94, 286, 458]. Product [446, 543, 381]. Products [435, 545, 474]. Program [400, 223]. Programming [15, 46, 295, 419, 453, 369, 273, 433, 268, 77, 391, 198, 494, 156, 241, 352, 500, 519, 355, 319, 536, 557, 376, 237, 329]. Programs [203, 264, 400, 2, 197, 105, 417]. Projection [156, 241]. Projective [352]. Proof [251, 500]. Propagated [201]. Proper [477]. Properties [320, 534, 179, 72, 263]. Property [301]. Pseudoinverses [344]. Punch [4]. Purely [524].

Quadratic [309, 453, 203, 275, 136, 391]. Quadrature [336, 207]. Quasi [337, 415]. Quasi-Dimensional [337]. Quasi-Linear [415]. Quasicyclic [83]. Quasilinearization [514]. Queue [540]. Queueing [29, 18]. Queues [351, 306].

R [256]. R. [555, 115]. Radar [165, 177]. Radiation [410, 311, 210, 307, 397, 191]. Radiative [498]. Radiometric [332]. Random [449, 148, 346, 230, 76, 261, 242, 497, 93, 285, 56, 508, 558, 470, 214, 296, 158, 485, 247]. Range [183]. Rank [118]. Rarefied [493]. Rate [551, 287]. Ratio [175]. Rational [362, 510, 39, 405, 504]. Reactor [100, 132]. Real [391, 547, 303, 457, 211, 504]. Realizability [290, 323]. Realizations [510]. Reciprocal [358]. Recognition [481]. Recollections [487]. Rectangular [515, 429]. Recurrence [475]. Recurring [108]. Recursion [414]. Reduction [113].

Redundancy [317, 466]. Reentrant [530]. Region [530]. Regression [261]. Regular [227]. Regularity [65]. Regulation [283, 316]. Related [379, 80, 526, 95]. Relation [44]. Relations [468]. Relativistic [258]. Relativity [313]. Relaxation [525]. Reliability [153, 288]. Remainder [318]. Remark [555, 256, 115]. Remarks [446, 395, 336, 39]. Rendezvous [236]. Replacement [34, 173, 176]. Representation [202, 220]. Representations [447, 25]. Research [1, 6, 12, 17, 21, 30]. Residue [534]. Resolving [329]. Resonance [300]. Resource [286]. Resources [5]. Respect [464]. Response [43, 49]. Restraining [76]. Restricted [93]. Result [185]. Results [86, 80]. Return [5]. Rev [91]. Review [97]. Revolution [495]. Riccati [516]. Rice [256]. Riemann [471]. Riemann-Like [471]. Ring [382, 452]. Ring-Stiffened [452]. Rings [534]. Rise [159]. Rods [544]. Root [74, 78, 251, 365, 366, 470]. Root-Locus [251]. Root-Squaring [366]. Roots [157, 135, 303, 211]. Rotating [495]. Rotations [79]. Round [398]. Ruin [527].

S [555]. S. [467]. Saddle [64]. Sample [136, 162, 239, 523]. Sampled [140]. Sampled-Data [140]. Sampling [346, 408, 158]. Satellites [385]. Scale [215, 81, 310]. Scaled [193]. Scattering [445, 64, 104, 16, 69]. Scheduling [52]. Scheme [271, 383]. Schemes [347, 380, 72, 154]. Schwinger [64, 472]. Screen [513, 307, 397]. Search [434, 138, 65, 348]. Seasonal [190]. Second [347, 549, 402, 200, 287, 291, 129]. Second-Degree [549]. Second-Order [347, 200, 287, 291]. Section [382]. Self [203, 433, 82]. Self-Adjoint [82]. Self-Dual [203]. Self-Duality [433]. Semi [226, 458, 60, 361]. Semi-Infinite [226, 361]. Semi-Iterative [60]. Semi-Markov [458].

Semigroups [499]. Sensitivity [332]. Sequence [160, 407]. Sequences [108]. Sequencing [268]. Sequential [113, 479, 65]. Series [41, 148, 226, 528, 167, 58, 373, 553]. Server [351, 540]. Service [351]. Set [350, 290, 323]. Sets [205, 372, 263]. Several [225, 208, 535, 480, 404, 149, 359]. Shallow [495, 90]. Shannon [456]. Shape [531]. Shapes [277]. Sharing [152]. Shear [436]. Shells [240, 410, 521, 428, 55, 495, 452, 90, 436]. Shock [531]. Short [500]. SIAM [238, 132, 467, 555, 282, 225, 439, 440, 556, 558, 270, 296, 397, 373, 239, 557]. Signal [339, 263]. Signals [40, 48, 424, 381, 522]. Significance [71]. Simple [82]. Simplification [130]. Simultaneous [215, 143, 454, 154, 552]. Single [29, 351, 540, 277]. Single-Server [540]. Singular [321, 411, 166, 164]. Singularities [45, 556, 220, 36]. Small [136]. Smoothing [15, 228, 66]. Solid [360]. Solids [358, 227]. Solution [412, 215, 289, 555, 226, 419, 143, 91, 265, 44, 454, 98, 390, 518, 398, 7, 450, 81, 354, 456, 39, 256, 367, 18, 26, 305, 276, 121]. Solutions [142, 151, 395, 478, 501, 302, 7, 104, 150, 272, 431, 120, 484, 292, 525, 287, 187, 342, 411, 438, 541, 109, 548]. Solving [471, 2, 28, 197, 42, 47, 425, 357]. Some [517, 20, 534, 406, 422, 505, 51, 477, 333, 67, 127, 302, 460, 18, 487, 404, 207, 548]. Sommerfeld [134]. Sorting [274]. Source [159, 531]. Space [462, 539, 305, 352]. Spaced [75]. Spaces [534, 391, 56, 475]. Spacing [162, 239, 523]. Specified [216]. Spectral [254, 136]. Spectrum [524]. Speed [111, 83]. Sphere [549, 360, 36]. Spherical [104, 521, 55, 90]. Spline [320]. Spontaneous [214, 296]. Springs [223]. Square [148, 398]. Squares [122, 343]. Squaring [366]. SSSR [91]. Stability [544, 161, 234, 347, 380, 44, 402, 537, 112, 438, 211]. Stabilization [512]. Stable [322, 271, 182]. Stage [20]. Stagewise [43, 49]. Standardization [130]. State [104, 360, 437]. Statement [124]. States [322, 521]. Stationary [64, 25, 549, 126, 540]. Statistical [192, 31, 9, 19, 23, 40, 48]. Steady [340, 104, 360]. Steady-State [104, 360]. Stefan [121]. Step [279, 155]. Stieltjes [185]. Stiffened [452]. Stochastic [190, 168, 458, 536, 557]. Stodola [284]. Straight [220]. Strategies [138]. Streamlines [189]. Strength [429]. Stress [235]. Stresses [55]. Strictly [243]. Strongly [326]. Structures [494]. Studies [297]. Study [173, 93]. Sturm [412, 284, 218]. Subject [317, 190]. Suboptimal [275]. Success [160]. Successive [465, 72, 154, 475, 60]. Sufficient [347, 380, 547]. Summation [41, 75, 221]. Sums [505, 209, 330, 331, 58, 373]. Supersonic [111, 491]. Supported [68]. Surface [230, 246, 429]. Surroundings [191]. Survey [67]. Switching [335, 456]. Symbols [217]. Symmetric [465, 490, 4, 298, 457, 531, 57, 452, 109, 468]. Symmetries [501]. Symmetry [301]. Synthesis [426]. System [91, 223, 541, 271, 383]. Systematic [434]. Systems [215, 234, 322, 503, 70, 29, 527, 463, 501, 72, 28, 204, 272, 450, 81, 509, 403, 466, 411, 130, 109, 184].

T [115]. Table [66]. Tables [248]. Tchebycheff [91, 116]. Tearing [333]. Technique [425, 172, 329]. Techniques [482, 212, 333, 267]. Telegrapher [292]. Temperature [410, 159, 259]. Temporal [35]. Tensor [543, 474]. Terminal [244]. Terminating [28, 167]. Test [162, 239]. Testing [427]. Tests [312]. th [188, 294]. their [117]. Theodore [487]. Theorem [453, 492, 358, 314, 360, 71, 342, 512, 500, 250, 519, 318]. Theorems [406, 408, 353, 392, 439, 526, 209]. Theories [128]. Theory [145, 20, 15, 33, 52, 364, 50, 140, 22, 350, 51, 228, 136, 358, 31, 9, 19, 23, 486, 493, 40, 48, 106, 494, 114, 330, 331, 313, 382, 47, 306, 386, 90, 3]. Thermal [410]. Thermo [357]. Thermo-Chemical [357]. Thermoelastic [360]. Thick [298]. Thickness [216]. Thin [24, 191, 216, 259]. Third [302]. Three [539, 137, 450]. Three-Dimensional [450]. Three-Point [137]. Time [316, 152, 528, 540, 310, 424, 413, 231, 236, 553]. Time-Invariant [231]. Time-Scale [310]. Time-Sharing [152]. Times [351]. Toeplitz [442]. Tool [24]. Topics [379]. Toroidal [436]. Torsion [4]. Traffic [377]. Trajectories [123]. Transcendental [256, 211]. Transfer [498, 191]. Transform [295, 448, 547]. Transformation [378, 371]. Transformations [534, 262, 110, 267]. Transforming [79]. Transforms [504]. Transient [43, 49]. Transition [437, 297]. Transitivity [372]. Transmission [463]. Transport [104, 150]. Transportation [488, 98, 400, 59, 257]. Transverse [449, 436]. Treating [218]. Treatment [252, 341, 257]. Trees [497]. Triangular [79]. Trichotomous [466]. Tridiagonal [113, 334, 440]. Tridiagonalization [550]. Truncation [148]. Turbulence [31, 9, 19, 23]. Turbulent [489, 297]. Turing [114]. Turning [370, 416, 130]. Turning-Point [370, 416]. Two [317, 327, 209, 514, 271, 383, 459]. Type [85, 284, 72, 302, 133, 166, 164, 383].

Uhlenbeck [281]. Unbiased [522]. Unbounded [127]. Unequally [75]. Unified [253, 341]. Uniform [263]. Uniqueness [323]. Unit [549]. Unitary [79]. Unknown [522]. Unstable [422, 324]. Upon [128]. Upper [227]. Use [73, 63, 284]. Used [333, 94]. Useful [110]. Using [289, 86, 175, 552, 60]. Utility [202].

Utilization [224].

V [21]. Value [406, 289, 347, 380, 321, 170, 360, 222, 270, 42, 459]. Valued [230, 261, 179]. Values [477, 549, 485, 319]. Varga [555]. Variable [284]. Variables [230, 539, 225, 208, 294, 535, 480, 404, 485, 149, 359, 247]. Variates [136]. Variation [151, 472]. Variational [46, 16, 255, 509, 47]. Variations [190]. Vector [534, 261, 179]. Vector-Valued [261, 179]. Velocities [432]. Versus [434]. Vertices [205, 290, 323, 409]. VI [30]. Vibrating [226, 341]. Vibration [223, 68]. Vibrations [449, 304, 235, 452]. View [40, 48]. Viscoelastic [358, 450]. Vortex [220]. Vortex-Like [220].

Waiting [540]. Walk [93, 285]. Walks [76, 56, 214, 296]. Wave [340]. Waveforms [119]. Waves [513, 531]. Weighted [209]. Which [381, 527, 201, 191]. Whittaker [408]. Whose [361]. Wielandt [101]. Wiener [238, 194, 228]. Wiener-Kolmogorov [228]. Without [467]. World [312]. WT [381]. WT-Product [381]. Wye [371]. Wye-Delta [371].

Zero [179, 191]. Zeros [443, 446]. Zuhovickiı [91].

References

Clippinger:1953:ADCa

[1] R. F. Clippinger, B. Dimsdale, and J. H. Levin. Automatic digital computers in industrial research. I. Journal of the Society for Industrial and Applied Mathematics, 1(1):1–15, September 1953. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hoffman:1953:CES

[2] A. Hoffman, M. Mannos, D. Sokolowsky, and N. Wiegmann. Computational experience in solving linear programs. Journal of the Society for Industrial and Applied Mathematics, 1(1):17–33, September 1953. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Zadeh:1953:TF

[3] L. A. Zadeh. Theory of filtering. Journal of the Society for Industrial and Applied Mathematics, 1(1):35–51, September 1953. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Payne:1953:ASP

[4] L. E. Payne. On axially symmetric punch, crack and torsion problems. Journal of the Society for Industrial and Applied Mathematics, 1(1):53–71, September 1953. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fite:1953:MRL

[5] Wade L. Fite. Maximization of return from limited resources. Journal of the Society for Industrial and Applied Mathematics, 1(2):73–90, December 1953. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Clippinger:1953:ADCb

[6] R. F. Clippinger, B. Dimsdale, and J. H. Levin. Automatic digital computers in industrial research, II. Journal of the Society for Industrial and Applied Mathematics, 1(2):91–110, December 1953. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Juncosa:1953:OCS

[7] M. L. Juncosa and D. M. Young. On the order of convergence of solutions of a difference equation to a solution of the diffusion equation. Journal of the Society for Industrial and Applied Mathematics, 1(2):111–135, December 1953. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Woodbury:1953:LCG

[8] Max A. Woodbury. Linear-convex games. Journal of the Society for Industrial and Applied Mathematics, 1(2): 137–142, December 1953. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

KampedeFeriet:1954:ISTa

[9] J. Kampe de Feriet and S. I. Pai. Introduction to the statistical theory of turbulence. I. Journal of the Society for Industrial and Applied Mathematics, 2 (1):1–9, March 1954. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Goldberg:1954:PMB

[10] Samuel Goldberg. Probability models in biology and engineering. Journal of the Society for Industrial and Applied Mathematics, 2(1):10–19, March 1954. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Abramowitz:1954:PEI

[11] Milton Abramowitz. On the practical evaluation of integrals. Journal of the Society for Industrial and Applied Mathematics, 2(1):20–35, March 1954. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Clippinger:1954:ADCa

[12] R. F. Clippinger, B. Dimsdale, and J. H. Levin. Automatic digital computers in industrial research. III. Journal of the Society for Industrial and Applied Mathematics, 2(1):36–56, March 1954. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Birkhoff:1954:CPD

[13] Garrett Birkhoff. Classification of partial differential equations. Journal of the Society for Industrial and Applied Mathematics, 2(1):57–67, March 1954. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gorn:1954:AAC

[14] Saul Gorn. The automatic analysis and control of computing errors. Journal of the Society for Industrial and Applied Mathematics, 2(2):69–81, June 1954. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bellman:1954:TDP

[15] R. Bellman, I. Glicksberg, and O. Gross. The theory of dynamic programming as applied to a smoothing problem. Journal of the Society for Industrial and Applied Mathematics, 2(2):82–88, June 1954. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kodis:1954:IVM

[16] Ralph D. Kodis. An introduction to variational methods in electromagnetic scattering. Journal of the Society for Industrial and Applied Mathematics, 2(2): 89–112, June 1954. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Clippinger:1954:ADCb

[17] R. F. Clippinger, B. Dimsdale, and J. H. Levin. Automatic digital computers in industrial research. IV. Journal of the Society for Industrial and Applied Mathematics, 2(2):113–131, June 1954. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

OBrien:1954:SSQ

[18] George G. O’Brien. The solution of some queueing problems. Journal of the Society for Industrial and Applied Mathematics, 2(3):133–142, September 1954. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

KampedeFeriet:1954:ISTb

[19] J. Kampe de Feriet and S. I. Pai. Introduction to the statistical theory of turbulence. II. Journal of the Society for Industrial and Applied Mathematics, 2 (3):143–174, September 1954. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bellman:1954:SCP

[20] Richard Bellman and Oliver Gross. Some combinatorial problems arising in the theory of multi-stage processes. Journal of the Society for Industrial and Applied Mathematics, 2(3):175–183, September 1954. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Clippinger:1954:ADCc

[21] R. F. Clippinger, B. Dimsdale, and J. H. Levin. Automatic digital computers in industrial research. V. Journal of the Society for Industrial and Applied Mathematics, 2(3):184–200, September 1954. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Burks:1954:CDN

[22] Arthur W. Burks, Robert McNaughton, Carl H. Pollmar, Don W. Warren, and Jesse B. Wright. Complete decoding nets: General theory and minimality. Journal of the Society for Industrial and Applied Mathematics, 2(4):201–243, December 1954. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

KampedeFeriet:1954:ISTc

[23] J. Kampe de Feriet. Introduction to the statistical theory of turbulence. III. Journal of the Society for Industrial and Applied Mathematics, 2(4):244–271, December 1954. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Loeb:1955:DCL

[24] Arthur L. Loeb and Harry H. Denman. The digital computer as a laboratory tool: Electronic digital computation of optical constants of thin metal films. Journal of the Society for Industrial and Applied Mathematics, 3(1):1–16, March 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Erdelyi:1955:ARF

[25] A. Erdélyi. Asymptotic representations of Fourier integrals and the method of stationary phase. Journal of the Society for Industrial and Applied Mathematics, 3(1):17–27, March 1955. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Peaceman:1955:NSP

[26] D. W. Peaceman and H. H. Rachford, Jr. The numerical solution of parabolic and elliptic differential equations. Journal of the Society for Industrial and Applied Mathematics, 3(1):28–41, March 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Douglas:1955:NII

[27] Jim Douglas, Jr. On the numerical integration \frac{∂^{2}u}{∂x^{2}} + \frac{∂^{2}u}{∂y^{2}} = \frac{∂u}{∂t} by implicit methods. Journal of the Society for Industrial and Applied Mathematics, 3(1): 42–65, March 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Householder:1955:TNI

[28] A. S. Householder. Terminating and nonterminating iterations for solving linear systems. Journal of the Society for Industrial and Applied Mathematics, 3 (2):67–72, June 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fagen:1955:QSS

[29] R. E. Fagen and John Riordan. Queueing systems for single and multiple operation. Journal of the Society for Industrial and Applied Mathematics, 3(2): 73–79, June 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Clippinger:1955:ADC

[30] R. F. Clippinger, B. Dimsdale, and J. H. Levin. Automatic digital computers in industrial research. VI. Journal of the Society for Industrial and Applied Mathematics, 3(2):80–89, June 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

KampeDeFeriet:1955:IST

[31] J. Kampe De Feriet and S. I. Pai. Introduction to the statistical theory of turbulence. IV. Journal of the Society for Industrial and Applied Mathematics, 3(2):90–117, June 1955. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

McMillan:1955:HP

[32] Brockway McMillan. History of a problem. Journal of the Society for Industrial and Applied Mathematics, 3(3): 119–128, September 1955. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bellman:1955:FEA

[33] Richard Bellman. A functional equation arising in allocation theory. Journal of the Society for Industrial and Applied Mathematics, 3(3):129–132, September 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bellman:1955:ERP

[34] Richard Bellman. Equipment replacement policy. Journal of the Society for Industrial and Applied Mathematics, 3 (3):133–136, September 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Putnam:1955:TMD

[35] C. R. Putnam. Temporal means and distribution functions. Journal of the Society for Industrial and Applied Mathematics, 3(3):137–141, September 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Yeh:1955:PDC

[36] G. C. K. Yeh, J. Martinek, and G. S. S. Ludford. The potentials due to certain singularities in the presence of a fixed sphere. Journal of the Society for Industrial and Applied Mathematics, 3(3): 142–152, September 1955. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cole:1955:ACN

[37] K. S. Cole, H. A. Antosiewicz, and P. Rabinowitz. Automatic computation of nerve excitation. Journal of the Society for Industrial and Applied Mathematics, 3(3):153–172, September 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Forsythe:1955:CCM

[38] George E. Forsythe. Computing constrained minima with Lagrange multipliers. Journal of the Society for Industrial and Applied Mathematics, 3(4): 173–178, December 1955. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Luke:1955:RMS

[39] Yudell L. Luke. Remarks on the τmethod for the solution of linear differential equations with rational coefficients. Journal of the Society for Industrial and Applied Mathematics, 3(4): 179–191, December 1955. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Middleton:1955:DES

[40] David Middleton and David van Meter. Detection and extraction of signals in noise from the point of view of statistical decision theory. I. Journal of the Society for Industrial and Applied Mathematics, 3(4):192–253, December 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Brachman:1955:NSS

[41] Malcolm K. Brachman. Notes on the summation of series. Journal of the Society for Industrial and Applied Mathematics, 3(4):254–258, December 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Reid:1955:MSC

[42] Walter P. Reid. A method for solving certain boundary value problems. Journal of the Society for Industrial and Applied Mathematics, 3(4):259–261, December 1955. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Acrivos:1956:TRSa

[43] Andreas Acrivos. The transient response of stagewise processes. Journal of the Society for Industrial and Applied Mathematics, 4(1):1–19, March 1956. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Douglas:1956:RBS

[44] Jim Douglas, Jr. On the relation between stability and convergence in the numerical solution of linear parabolic and hyperbolic differential equations. Journal of the Society for Industrial and Applied Mathematics, 4(1):20–37, March 1956. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Erdelyi:1956:AEF

[45] A. Erdélyi. Asymptotic expansions of Fourier integrals involving logarithmic singularities. Journal of the Society for Industrial and Applied Mathematics, 4 (1):38–47, March 1956. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See note [556].

Bellman:1956:VPC

[46] Richard Bellman. A variational problem with constraints in dynamic programming. Journal of the Society for Industrial and Applied Mathematics, 4(1): 48–61, March 1956. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Stiefel:1956:SFI

[47] E. Stiefel. On solving Fredholm integral equations: Applications to conformal mapping and variational problems of potential theory. Journal of the Society for Industrial and Applied Mathematics, 4(2):63–85, June 1956. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Middleton:1956:DES

[48] David Middleton and David van Meter. Detection and extraction of signals in noise from the point of view of statistical decision theory. II. Journal of the Society for Industrial and Applied Mathematics, 4(2):86–119, June 1956. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Acrivos:1956:TRSb

[49] Andreas Acrivos. The transient response of stagewise processes. Journal of the Society for Industrial and Applied Mathematics, 4(2):120–130, June 1956. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bendat:1956:GTL

[50] Julius Bendat. A general theory of linear prediction and filtering. Journal of the Society for Industrial and Applied Mathematics, 4(3):131–151, September 1956. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cohn:1956:SAN

[51] Harvey Cohn. Some applied number theory. Journal of the Society for Industrial and Applied Mathematics, 4(3): 152–167, September 1956. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bellman:1956:MAS

[52] Richard Bellman. Mathematical aspects of scheduling theory. Journal of the Society for Industrial and Applied Mathematics, 4(3):168–205, September 1956. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Stegun:1956:PC

[53] Irene A. Stegun and Milton Abramowitz. Pitfalls in computation. Journal of the Society for Industrial and Applied Mathematics, 4(4):207–219, December 1956. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Arms:1956:MBI

[54] R. J. Arms, L. D. Gates, and B. Zondek. A method of block iteration. Journal of the Society for Industrial and Applied Mathematics, 4(4):220–229, December 1956. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Reissner:1956:NMB

[55] Eric Reissner. A note on membrane and bending stresses in spherical shells. Journal of the Society for Industrial and Applied Mathematics, 4(4):230–240, December 1956. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Montroll:1956:RWM

[56] Elliot W. Montroll. Random walks in multidimensional spaces, especially on periodic lattices. Journal of the Society for Industrial and Applied Mathematics, 4(4):241–260, December 1956. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Schlesinger:1957:AEE

[57] Stewart Schlesinger. Approximating eigenvalues and eigenfunctions of symmetric kernels. Journal of the Society for Industrial and Applied Mathematics, 5(1):1–14, March 1957. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Vinti:1957:SCS

[58] John P. Vinti and Tadeusz Leser. The sums of certain series involving Bessel functions. Journal of the Society for Industrial and Applied Mathematics, 5(1): 15–31, March 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See errata [373].

Munkres:1957:AAT

[59] James Munkres. Algorithms for the assignment and transportation problems. Journal of the Society for Industrial and Applied Mathematics, 5(1):32–38, March 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Varga:1957:CSO

[60] Richard S. Varga. A comparison of the successive overrelaxation method and semi-iterative methods using Chebyshev polynomials. Journal of the Society for Industrial and Applied Mathematics, 5 (2):39–46, June 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Krzywoblocki:1957:GBL

[61] M. Z. V. Krzywoblocki and H. A. Hassan. Generalization of Bergman’s linear integral operator method to diabatic flow. Journal of the Society for Industrial and Applied Mathematics, 5(2): 47–65, June 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Peirce:1957:NIP

[62] William H. Peirce. Numerical integration over the planar annulus. Journal of the Society for Industrial and Applied Mathematics, 5(2):66–73, June 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Forsythe:1957:GUO

[63] George E. Forsythe. Generation and use of orthogonal polynomials for datafitting with a digital computer. Journal of the Society for Industrial and Applied Mathematics, 5(2):74–88, June 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Dolph:1957:SPC

[64] C. L. Dolph. A saddle point characterization of the Schwinger stationary points in exterior scattering problems. Journal of the Society for Industrial and Applied Mathematics, 5(3):89– 104, September 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kiefer:1957:OSS

[65] J. Kiefer. Optimum sequential search and approximation methods under minimum regularity assumptions. Journal of the Society for Industrial and Applied Mathematics, 5(3):105–136, September 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Greville:1957:SFT

[66] T. N. E. Greville. On smoothing a finite table: A matrix approach. Journal of the Society for Industrial and Applied Mathematics, 5(3):137–154, September 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Householder:1957:SSC

[67] Alston S. Householder. A survey of some closed methods for inverting matrices. Journal of the Society for Industrial and Applied Mathematics, 5(3): 155–169, September 1957. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Payne:1957:LBV

[68] L. E. Payne and H. F. Weinberger. Lower bounds for vibration frequencies of elastically supported membranes and plates. Journal of the Society for Industrial and Applied Mathematics, 5(4):

171–182, December 1957. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sims:1957:CAI

[69] A. R. Sims. Certain aspects of the inverse scattering problem. Journal of the Society for Industrial and Applied Mathematics, 5(4):183–205, December 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Brown:1957:ADL

[70] WM. M. Brown. Analysis of discrete linear systems. Journal of the Society for Industrial and Applied Mathematics, 5 (4):206–224, December 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Riordan:1957:CST

[71] John Riordan. The combinatorial significance of a theorem of Polya. Journal of the Society for Industrial and Applied Mathematics, 5(4):225–237, December 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Heller:1957:OPL

[72] J. Heller. Ordering properties of linear successive iteration schemes applied to multi-diagonal type linear systems. Journal of the Society for Industrial and Applied Mathematics, 5(4):238–243, December 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Campbell:1957:UHE

[73] L. Lorne Campbell. On the use of Hermite expansions in noise problems. Journal of the Society for Industrial and Applied Mathematics, 5(4):244–249, December 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Brauer:1957:MCG

[74] Alfred Brauer. A method for the computation of the greatest root of a positive matrix. Journal of the Society for Industrial and Applied Mathematics, 5 (4):250–253, December 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Salzer:1957:NMI

[75] Herbert E. Salzer. Note on multivariate interpolation for unequally spaced arguments, with an application to double summation. Journal of the Society for Industrial and Applied Mathematics, 5(4):254–262, December 1957. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fagen:1958:RWR

[76] R. E. Fagen and T. A. Lehrer. Random walks with restraining barrier as applied to the biased binary counter. Journal of the Society for Industrial and Applied Mathematics, 6(1):1–14, March 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kelley:1958:ALP

[77] James E. Kelley, Jr. An application of linear programming to curve fitting. Journal of the Society for Industrial and Applied Mathematics, 6(1):15–22, March 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Dimsdale:1958:NCC

[78] B. Dimsdale. The non-convergence of a characteristic root method. Journal of the Society for Industrial and Applied Mathematics, 6(1):23–25, March 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Givens:1958:CPU

[79] Wallace Givens. Computation of plane unitary rotations transforming a general matrix to triangular form. Journal of the Society for Industrial and Applied Mathematics, 6(1):26–50, March 1958. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hestenes:1958:IMB

[80] Magnus R. Hestenes. Inversion of matrices by biorthogonalization and related results. Journal of the Society for Industrial and Applied Mathematics, 6(1): 51–90, March 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lanczos:1958:ISL

[81] C. Lanczos. Iterative solution of largescale linear systems. Journal of the Society for Industrial and Applied Mathematics, 6(1):91–109, March 1958. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wittmeyer:1958:NMD

[82] H. Wittmeyer. A new method for developing simple formulae for the eigenvalues of linear ordinary self-adjoint differential equations. Journal of the Society for Industrial and Applied Mathematics, 6(2):111–143, June 1958. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Henrici:1958:SCC

[83] Peter Henrici. On the speed of convergence of cyclic and quasicyclic Jacobi methods for computing eigenvalues of Hermitian matrices. Journal of the Society for Industrial and Applied Mathematics, 6(2):144–162, June 1958. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Tarnove:1958:DEM

[84] Ivin Tarnove. Determination of eigenvalues of matrices having polynomial elements. Journal of the Society for Industrial and Applied Mathematics, 6(2): 163–171, June 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Causey:1958:CEN

[85] Robert L. Causey. Computing eigenvalues of non-Hermitian matrices by methods of Jacobi type. Journal of the Society for Industrial and Applied Mathematics, 6(2):172–181, June 1958. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gregory:1958:RUL

[86] Robert T. Gregory. Results using Lanczos’ method for finding eigenvalues of arbitrary matrices. Journal of the Society for Industrial and Applied Mathematics, 6(2):182–188, June 1958. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Householder:1958:CMI

[87] Alston S. Householder. A class of methods for inverting matrices. Journal of the Society for Industrial and Applied Mathematics, 6(2):189–195, June 1958. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cole:1958:ACN

[88] K. S. Cole, H. A. Antosiewicz, and P. Rabinowitz. Automatic computation of nerve excitation, correction. Journal of the Society for Industrial and Applied Mathematics, 6(2):196–197, June 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kochen:1958:MFI

[89] Manfred Kochen. A mathematical formulation of influence distributions in decision-making groups. Journal of the Society for Industrial and Applied Mathematics, 6(3):199–208, September 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Weinitschke:1958:NTS

[90] Hubertus J. Weinitschke. On the nonlinear theory of shallow spherical shells. Journal of the Society for Industrial and Applied Mathematics, 6(3):209–232, September 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cheney:1958:NPZ

[91] Ward Cheney and Allen A. Goldstein. Note on a paper by Zuhovickiĭ concerning the Tchebycheff problem for linear equations: [“An Algorithm for the Solution of the Çebyšev Approximation Problem in the Case of a Finite System of Incompatible Linear Equations”, Doklady Akad. Nauk SSSR (N.S.), 79 (1951), pp, 561–564; Math. Rev. 13 (1952), p. 285]. Journal of the Society for Industrial and Applied Mathematics, 6(3):233–239, September 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bridgland:1958:NNI

[92] T. F. Bridgland, Jr. A note on numerical integrating operators. Journal of the Society for Industrial and Applied Mathematics, 6(3):240–256, September 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lehman:1958:SRR

[93] R. Sherman Lehman and George H. Weiss. A study of the restricted random walk. Journal of the Society for Industrial and Applied Mathematics, 6 (3):257–278, September 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Osborne:1958:AMD

[94] Elmer E. Osborne. On acceleration and matrix deflation processes used with the power method. Journal of the Society for Industrial and Applied Mathematics, 6(3):279–287, September 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Thacher:1958:GCR

[95] Henry C. Thacher, Jr. Generalization of concepts related to linear dependence. Journal of the Society for Industrial and Applied Mathematics, 6(3): 288–299, September 1958. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Whaples:1958:NDI

[96] G. Whaples. A note on degree-n independence. Journal of the Society for Industrial and Applied Mathematics, 6 (3):300–301, September 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Maradudin:1958:DLP

[97] Alexei Maradudin and George H. Weiss. The disordered lattice problem: A review. Journal of the Society for Industrial and Applied Mathematics, 6(3): 302–319, September 1958. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gerstenhaber:1958:SMT

[98] Murray Gerstenhaber. A solution method for the transportation problem. Journal of the Society for Industrial and Applied Mathematics, 6(4): 321–334, December 1958. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Guderley:1958:NEP

[99] K. G. Guderley. On nonlinear eigenvalue problems for matrices. Journal of the Society for Industrial and Applied Mathematics, 6(4):335–353, December 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Birkhoff:1958:RCN

[100] Garrett Birkhoff and Richard S. Varga.

Reactor criticality and nonnegative matrices. Journal of the Society for Industrial and Applied Mathematics, 6(4): 354–377, December 1958. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See errata [132].

Frank:1958:CEC

[101] Werner L. Frank. Computing eigenvalues of complex matrices by determinant evaluation and by methods of Danilewski and Wielandt. Journal of the Society for Industrial and Applied Mathematics, 6(4):378–392, December 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

White:1958:CEE

[102] Paul A. White. The computation of eigenvalues and eigenvectors of a matrix. Journal of the Society for Industrial and Applied Mathematics, 6(4): 393–437, December 1958. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bauer:1958:MCM

[103] W. F. Bauer. The Monte Carlo method. Journal of the Society for Industrial and Applied Mathematics, 6(4):438–451, December 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Keller:1958:AST

[104] Herbert B. Keller. Approximate solutions of transport problems — part I: Steady-state, elastic scattering in plane and spherical geometry. Journal of the Society for Industrial and Applied Mathematics, 6(4):452–465, December 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Newman:1958:EMI

[105] Morris Newman and John Todd. The evaluation of matrix inversion programs. Journal of the Society for Industrial and Applied Mathematics, 6(4):466–476, December 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Pawlak:1959:DNT

[106] Zdzislaw Pawlak. Decoding nets and the theory of graphs. Journal of the Society for Industrial and Applied Mathematics, 7(1):1–5, March 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Leipnik:1959:IEB

[107] Roy Leipnik. Integral equations, biorthonormal expansions, and noise. Journal of the Society for Industrial and Applied Mathematics, 7(1):6–30, March 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Zierler:1959:LRS

[108] Neal Zierler. Linear recurring sequences.

Journal of the Society for Industrial and Applied Mathematics, 7(1):31–48, March 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Weinberger:1959:EBF

[109] H. F. Weinberger. Error bounds in finite-difference approximation to solutions of symmetric hyperbolic systems. Journal of the Society for Industrial and Applied Mathematics, 7(1):49–75, March 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Grosswald:1959:TUN

[110] Emil Grosswald. Transformations useful in numerical integration methods. Journal of the Society for Industrial and Applied Mathematics, 7(1):76–84, March 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ehlers:1959:MCI

[111] F. Edward Ehlers. The method of characteristics for isoenergetic supersonic flows adapted to high-speed digital computers. Journal of the Society for Industrial and Applied Mathematics, 7(1): 85–100, March 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Putnam:1959:SIH

[112] C. R. Putnam. On the stability intervals of the Hill equation. Journal of the Society for Industrial and Applied Mathematics, 7(1):101–106, March 1959. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bauer:1959:SRT

[113] F. L. Bauer. Sequential reduction to tridiagonal form. Journal of the Society for Industrial and Applied Mathematics, 7(1):107–113, March 1959. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rogers:1959:PTT

[114] Hartley Rogers, Jr. The present theory of Turing machine computability. Journal of the Society for Industrial and Applied Mathematics, 7(1):114–130, March 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Weisfeld:1959:RPR

[115] Morris Weisfeld. A remark on a paper of R. T. Gregory. Journal of the Society for Industrial and Applied Mathematics, 7 (1):131, March 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rice:1959:CAB

[116] J. R. Rice. On the convergence of an algorithm for best Tchebycheff approximations. Journal of the Society for Industrial and Applied Mathematics, 7(2): 133–142, June 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Baker:1959:NDN

[117] George A. Baker, Jr. A new derivation of Newton’s identities and their application to the calculation of the eigenvalues of a matrix. Journal of the Society for Industrial and Applied Mathematics, 7(2): 143–148, June 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wilf:1959:MIA

[118] Herbert S. Wilf. Matrix inversion by the annihilation of rank. Journal of the Society for Industrial and Applied Mathematics, 7(2):149–151, June 1959. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Swerling:1959:PEW

[119] P. Swerling. Parameter estimation for waveforms in additive Gaussian noise. Journal of the Society for Industrial and Applied Mathematics, 7(2):152– 166, June 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lees:1959:ASP

[120] Milton Lees. Approximate solutions of parabolic equations. Journal of the Society for Industrial and Applied Mathematics, 7(2):167–183, June 1959. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Trench:1959:EMS

[121] William F. Trench. On an explicit method for the solution of a Stefan problem. Journal of the Society for Industrial and Applied Mathematics, 7(2): 184–204, June 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Blum:1959:BEL

[122] Marvin Blum and Hirsch Gottschalk. On the bias errors of least squares digital filters. Journal of the Society for Industrial and Applied Mathematics, 7(2): 205–214, June 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Breakwell:1959:OT

[123] John V. Breakwell. The optimization of trajectories. Journal of the Society for Industrial and Applied Mathematics, 7 (2):215–247, June 1959. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Erdelyi:1959:IS

[124] A. Erdélyi. Introductory statement. Journal of the Society for Industrial and Applied Mathematics, 7(3):249–252, September 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

vanderCorput:1959:N

[125] J. G. van der Corput. Neutrices. Journal of the Society for Industrial and Applied Mathematics, 7(3):253–279, September 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Friedman:1959:SPN

[126] B. Friedman. Stationary phase with neighboring critical points. Journal of the Society for Industrial and Applied Mathematics, 7(3):280–289, September 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hull:1959:SIE

[127] T. E. Hull. On some integral equations with unbounded eigenvalues. Journal of the Society for Industrial and Applied Mathematics, 7(3):290–297, September 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Langer:1959:ATL

[128] Rudolph E. Langer. Asymptotic theories for linear ordinary differential equations depending upon a parameter. Journal of the Society for Industrial and Applied Mathematics, 7(3):298–305, September 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Olver:1959:LDE

[129] F. W. J. Olver. Linear differential equations of the second order with a large parameter. Journal of the Society for Industrial and Applied Mathematics, 7 (3):306–310, September 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Turrittin:1959:SSS

[130] H. L. Turrittin and W. A. Harris, Jr.

Standardization and simplification of systems of linear differential equations involving a turning point. Journal of the Society for Industrial and Applied Mathematics, 7(3):316–324, September 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wyman:1959:ABL

[131] Max Wyman. The asymptotic behavior of the Laurent coefficients. Journal of the Society for Industrial and Applied Mathematics, 7(3):325–342, September 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Birkhoff:1959:ERC

[132] Garrett Birkhoff and Richard S. Varga. Errata: “Reactor Criticality and Nonnegative Matrices” [J. SIAM 6 (1958), no. 4, 354–377]. Journal of the Society for Industrial and Applied Mathematics, 7(3):343, September 1959. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [100].

Salzer:1959:BAM

[133] Herbert E. Salzer. Best approximation of mixed type. Journal of the Society for Industrial and Applied Mathematics, 7(4):345–360, December 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Conte:1959:NIO

[134] S. D. Conte and J. W. Miles. On the numerical integration of the Orr– Sommerfeld equation. Journal of the Society for Industrial and Applied Mathematics, 7(4):361–366, December 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Flanagan:1959:ERC

[135] Czerna Flanagan and J. E. Maxfield. Estimates of the roots of certain polynomials. Journal of the Society for Industrial and Applied Mathematics, 7(4): 367–373, December 1959. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Grenander:1959:DQF

[136] ULF Grenander, H. O. Pollak, and D. Slepian. The distribution of quadratic forms in normal variates: A small sample theory with applications to spectral analysis. Journal of the Society for Industrial and Applied Mathematics, 7(4): 374–401, December 1959. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hull:1959:EBF

[137] T. E. Hull and A. C. R. Newbery. Error bounds for a family of three-point integration procedures. Journal of the Society for Industrial and Applied Mathematics, 7(4):402–412, December 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gilbert:1959:OSS

[138] E. N. Gilbert. Optimal search strategies. Journal of the Society for Industrial and Applied Mathematics, 7(4): 413–424, December 1959. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Radok:1959:MFE

[139] J. R. M. Radok. Method of functional extrapolation for the numerical integration of ordinary differential equations. Journal of the Society for Industrial and Applied Mathematics, 7(4): 425–430, December 1959. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bridgland:1959:LAF

[140] T. F. Bridgland, Jr. A linear algebraic formulation of the theory of sampleddata control. Journal of the Society for Industrial and Applied Mathematics, 7 (4):431–446, December 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fitzhugh:1959:ACN

[141] R. Fitzhugh and H. A. Antosiewicz. Automatic computation of nerve excitation — detailed corrections and additions. Journal of the Society for Industrial and Applied Mathematics, 7(4): 447–458, December 1959. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ablow:1959:ISD

[142] C. M. Ablow and C. L. Perry. Iterative solutions of the Dirichlet problem for Δu = u^{2}. Journal of the Society for Industrial and Applied Mathematics, 7 (4):459–467, December 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Chang:1959:NSS

[143] K. T. Chang. On the numerical solution of simultaneous linear differential equations. Journal of the Society for Industrial and Applied Mathematics, 7 (4):468–472, December 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Boyer:1959:IDE

[144] R. H. Boyer. An integro-differential equation for a Markov process. Journal of the Society for Industrial and Applied Mathematics, 7(4):473–486, December 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Akers:1959:TBF

[145] Sheldon B. Akers, Jr. On a theory of Boolean functions. Journal of the Society for Industrial and Applied Mathematics, 7(4):487–498, December 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bauer:1960:A

[146] Friedrich L. Bauer. The g-algorithm.

Journal of the Society for Industrial and Applied Mathematics, 8(1):1–17, March 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Salzer:1960:HGO

[147] Herbert E. Salzer. Hermite’s general osculatory interpolation formula and a finite difference analogue. Journal of the Society for Industrial and Applied Mathematics, 8(1):18–27, March 1960. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Brown:1960:MST

[148] J. L. Brown, Jr. Mean square truncation error in series expansions of random functions. Journal of the Society for Industrial and Applied Mathematics, 8(1): 28–32, March 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Thacher:1960:ISV

[149] Henry C. Thacher, Jr. and W. E. Milne.

Interpolation in several variables. Journal of the Society for Industrial and Applied Mathematics, 8(1):33–42, March 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Keller:1960:AST

[150] Herbert B. Keller. Approximate solutions of transport problems. II. convergence and applications of the discrete ordinate method. Journal of the Society for Industrial and Applied Mathematics, 8(1):43–73, March 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Blair:1960:PSU

[151] K. W. Blair and W. S. Loud. Periodic solutions of x′′ + cx′ + g(x) = Ef(t) under variation of certain parameters. Journal of the Society for Industrial and Applied Mathematics, 8(1):74–101, March 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Boyell:1960:ATS

[152] Roger L. Boyell. Analysis of timesharing in digital computers. Journal of the Society for Industrial and Applied Mathematics, 8(1):102–124, March 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Drenick:1960:MAR

[153] R. F. Drenick. Mathematical aspects of the reliability problem. Journal of the Society for Industrial and Applied Mathematics, 8(1):125–149, March 1960. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Heller:1960:SSA

[154] J. Heller. Simultaneous, successive and alternating direction iteration schemes. Journal of the Society for Industrial and Applied Mathematics, 8(1):150– 173, March 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Tauber:1960:DEC

[155] Selmo Tauber and Donald Dean. On difference equations containing step and delta functions. Journal of the Society for Industrial and Applied Mathematics, 8(1):174–180, March 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rosen:1960:GPM

[156] J. B. Rosen. The gradient projection method for nonlinear programming. part I. linear constraints. Journal of the Society for Industrial and Applied Mathematics, 8(1):181–217, March 1960. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Dimsdale:1960:CRM

[157] B. Dimsdale. Characteristic roots of a matric polynomial. Journal of the Society for Industrial and Applied Mathematics, 8(1):218–223, March 1960. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Shapiro:1960:AFS

[158] Harold S. Shapiro and Richard A. Silverman. Alias-free sampling of random noise. Journal of the Society for Industrial and Applied Mathematics, 8(2): 225–248, June 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lewis:1960:TRI

[159] J. A. Lewis and T. D. Riney. Temperature rise in an infinite medium heated by a planar source. Journal of the Society for Industrial and Applied Mathematics, 8(2):249–271, June 1960. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Anselone:1960:PES

[160] Philip M. Anselone. Persistence of an effect of a success in a Bernoulli sequence. Journal of the Society for Industrial and Applied Mathematics, 8(2): 272–279, June 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Arrow:1960:SGP

[161] Kenneth J. Arrow and Leonid Hurwicz. Stability of the gradient process in nperson games. Journal of the Society for Industrial and Applied Mathematics, 8 (2):280–294, June 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Weiss:1960:TFB

[162] Lionel Weiss. A test of fit based on the largest sample spacing. Journal of the Society for Industrial and Applied Mathematics, 8(2):295–299, June 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See erratum [239].

Reed:1960:PCC

[163] I. S. Reed and G. Solomon. Polynomial codes over certain finite fields. Journal of the Society for Industrial and Applied Mathematics, 8(2):300–304, June 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Struble:1960:DSI

[164] Raimond A. Struble. Differentiation of singular integrals of Cauchy type. Journal of the Society for Industrial and Applied Mathematics, 8(2):305–308, June 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kelly:1960:DREa

[165] E. J. Kelly, I. S. Reed, and W. L.

Root. The detection of radar echoes in noise. Journal of the Society for Industrial and Applied Mathematics, 8(2): 309–341, June 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Stewart:1960:NES

[166] Charles E. Stewart. On the numerical evaluation of singular integrals of Cauchy type. Journal of the Society for Industrial and Applied Mathematics, 8 (2):342–353, June 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Roe:1960:AMT

[167] Glenn M. Roe. An accurate method for terminating asymptotic series. Journal of the Society for Industrial and Applied Mathematics, 8(2):354–367, June 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Katz:1960:SME

[168] Stanley Katz. A stochastic model for emulsion polymerization. Journal of the Society for Industrial and Applied Mathematics, 8(2):368–375, June 1960. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Parter:1960:EEC

[169] S. Parter. On the eigenvalues and eigenvectors of a class of matrices. Journal of the Society for Industrial and Applied Mathematics, 8(2):376–388, June 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Johnson:1960:PEC

[170] M. W. Johnson and E. Reissner. Parametric expansions for a class of boundary value problems of partial differential equations. Journal of the Society for Industrial and Applied Mathematics, 8(2): 389–396, June 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Mills:1960:EPF

[171] Harlan Mills. Equilibrium points in finite games. Journal of the Society for Industrial and Applied Mathematics, 8(2): 397–402, June 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wachspress:1960:ADI

[172] E. L. Wachspress and G. J. Habetler. An alternating-direction-implicit iteration technique. Journal of the Society for Industrial and Applied Mathematics, 8(2):403–424, June 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Dreyfus:1960:GER

[173] Stuart E. Dreyfus. A generalized equipment replacement study. Journal of the Society for Industrial and Applied Mathematics, 8(3):425–435, September 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

McKinney:1960:MED

[174] Wilbur T. McKinney. Matrix equations for the determination of beam natural frequencies. Journal of the Society for Industrial and Applied Mathematics, 8 (3):436–457, September 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Loeb:1960:ACA

[175] Henry L. Loeb. Algorithms for Chebyshev approximations using the ratio of linear forms. Journal of the Society for Industrial and Applied Mathematics, 8 (3):458–465, September 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sinden:1960:RED

[176] Frank W. Sinden. The replacement and expansion of durable equipment. Journal of the Society for Industrial and Applied Mathematics, 8(3):466–480, September 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kelly:1960:DREb

[177] E. J. Kelly, I. S. Reed, and W. L.

Root. The detection of radar echoes in noise. Journal of the Society for Industrial and Applied Mathematics, 8(3): 481–507, September 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wendroff:1960:CDO

[178] Burton Wendroff. On the convergence of the discrete ordinate method. Journal of the Society for Industrial and Applied Mathematics, 8(3):508–513, September 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gerstenhaber:1960:PFZ

[179] M. Gerstenhaber. A procedure for finding a zero of a vector-valued function with certain monotonicity properties. Journal of the Society for Industrial and Applied Mathematics, 8(3):

514–517, September 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lees:1960:GP

[180] Milton Lees. The Goursat problem. Journal of the Society for Industrial and Applied Mathematics, 8(3):518–530, September 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bridgland:1960:NNI

[181] T. F. Bridgland, Jr. A note on numerical integrating operators. II. Journal of the Society for Industrial and Applied Mathematics, 8(3):531–536, September 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wilf:1960:MSN

[182] Herbert S. Wilf. Maximally stable numerical integration. Journal of the Society for Industrial and Applied Mathematics, 8(3):537–540, September 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Franklin:1960:RFA

[183] J. N. Franklin. The range of a fleet of aircraft. Journal of the Society for Industrial and Applied Mathematics, 8(3): 541–548, September 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wendroff:1960:CDE

[184] Burton Wendroff. On centered difference equations for hyperbolic systems. Journal of the Society for Industrial and Applied Mathematics, 8(3):549–555, September 1960. CODEN JSIMAV.

ISSN 0368-4245 (print), 2168-3484 (electronic).

Moore:1960:EER

[185] John T. Moore. An extension of an early result of stieltjes. Journal of the Society for Industrial and Applied Mathematics, 8(4):557–559, December 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Keller:1960:PCD

[186] Herbert B. Keller. On the pointwise convergence of the discrete-ordinate method. Journal of the Society for Industrial and Applied Mathematics, 8(4): 560–567, December 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rauch:1960:ISA

[187] L. M. Rauch and W. C. Riddell. The iterative solutions of the analytical Nbody problem. Journal of the Society for Industrial and Applied Mathematics, 8(4):568–581, December 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bellman:1960:BP

[188] Richard Bellman and Robert Kalaba.

On k th best policies. Journal of the Society for Industrial and Applied Mathematics, 8(4):582–588, December 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sedney:1960:NDS

[189] R. Sedney, N. Gerber, and J. M. Bartos.

Numerical determination of streamlines from density data. Journal of the Society for Industrial and Applied Mathematics, 8(4):589–610, December 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Karlin:1960:OPD

[190] Samuel Karlin. Optimal policy for dynamic inventory process with stochastic demands subject to seasonal variations. Journal of the Society for Industrial and Applied Mathematics, 8(4): 611–629, December 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wilkins:1960:MMT

[191] J. Ernest Wilkins, Jr. Minimum-mass thin fins which transfer heat only by radiation to surroundings at absolute zero. Journal of the Society for Industrial and Applied Mathematics, 8(4): 630–639, December 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Albert:1960:SMP

[192] G. E. Albert. Statistical methods in prediction, filtering, and detection problems. Journal of the Society for Industrial and Applied Mathematics, 8(4): 640–653, December 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kellogg:1960:SIL

[193] R. Bruce Kellogg and L. C. Noderer. Scaled iterations and linear equations. Journal of the Society for Industrial and Applied Mathematics, 8(4):654–661, December 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Beutler:1960:AWO

[194] Frederick J. Beutler. Approximations to Wiener optimum filters and predictors. Journal of the Society for Industrial and Applied Mathematics, 8(4): 662–679, December 1960. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See errata [238].

Drenick:1960:FLC

[195] R. F. Drenick. The failure law of complex equipment. Journal of the Society for Industrial and Applied Mathematics, 8(4):680–690, December 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rice:1960:CA

[196] John R. Rice. Chebyshev approximation by ^{∗}ab^{x} + c. Journal of the Society for Industrial and Applied Mathematics, 8 (4):691–702, December 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kelley:1960:CPM

[197] J. E. Kelley, Jr. The cutting-plane method for solving convex programs. Journal of the Society for Industrial and Applied Mathematics, 8(4):703–712, December 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lemke:1961:CGM

[198] C. E. Lemke. The constrained gradient method of linear programming. Journal of the Society for Industrial and Applied Mathematics, 9(1):1–17, March 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fulkerson:1961:KMM

[199] D. R. Fulkerson. An out-of-kilter method for minimal-cost flow problems. Journal of the Society for Industrial and Applied Mathematics, 9(1):18–27, March 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Morrison:1961:DEP

[200] D. D. Morrison. The deflation of eigenvalue problems with second-order ordinary differential operators. Journal of the Society for Industrial and Applied Mathematics, 9(1):28–30, March 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hull:1961:IPW

[201] T. E. Hull and A. C. R. Newbery. Integration procedures which minimize propagated errors. Journal of the Society for Industrial and Applied Mathematics, 9(1):31–47, March 1961. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fleischer:1961:NRU

[202] Isidore Fleischer. Numerical representation of utility. Journal of the Society for Industrial and Applied Mathematics, 9 (1):48–50, March 1961. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Dorn:1961:SDQ

[203] W. S. Dorn. Self-dual quadratic programs. Journal of the Society for Industrial and Applied Mathematics, 9(1): 51–54, March 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Jewell:1961:OSE

[204] James G. Jewell. Ordering systems of equations. Journal of the Society for Industrial and Applied Mathematics, 9(1): 55–71, March 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Balinski:1961:AFA

[205] M. L. Balinski. An algorithm for finding all vertices of convex polyhedral sets. Journal of the Society for Industrial and Applied Mathematics, 9(1):72–88, March 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Householder:1961:DM

[206] A. S. Householder. On deflating matrices. Journal of the Society for Industrial and Applied Mathematics, 9(1): 89–93, March 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Squire:1961:SAQ

[207] William Squire. Some applications of quadrature by differentiation. Journal of the Society for Industrial and Applied Mathematics, 9(1):94–108, March 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Greville:1961:NFF

[208] T. N. E. Greville. Note on fitting of functions of several independent variables. Journal of the Society for Industrial and Applied Mathematics, 9(1): 109–115, March 1961. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See erratum [225].

Myers:1961:TTM

[209] B. R. Myers. Two theorems in multiweighted sums. Journal of the Society for Industrial and Applied Mathematics, 9(1):116–126, March 1961. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rouse:1961:MNC

[210] Carl A. Rouse. A method for the numerical calculation of hydrodynamic flow and radiation diffusion by implicit differencing. Journal of the Society for Industrial and Applied Mathematics, 9 (1):127–135, March 1961. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wright:1961:SCR

[211] E. M. Wright. Stability criteria and the real roots of a transcendental equation. Journal of the Society for Industrial and Applied Mathematics, 9(1): 136–148, March 1961. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ehrlich:1961:FDT

[212] Louis W. Ehrlich, James D. Riley, W. Gilbert Strang, and B. Andreas Troesch. Finite-difference techniques for a boundary problem with an eigenvalue in a boundary condition. Journal of the Society for Industrial and Applied Mathematics, 9(1):149–164, March 1961. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Howland:1961:GCF

[213] James Lucien Howland and Remi Vaillancourt. A generalized curve-fitting procedure. Journal of the Society for Industrial and Applied Mathematics, 9(2): 165–168, June 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rosenstock:1961:RWS

[214] Herbert B. Rosenstock. Random walks with spontaneous emission. Journal of the Society for Industrial and Applied Mathematics, 9(2):169–188, June 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See errata [296].

Bellar:1961:ISL

[215] Fred J. Bellar, Jr. An iterative solution of large-scale systems of simultaneous linear equations. Journal of the Society for Industrial and Applied Mathematics, 9(2):189–193, June 1961. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wilkins:1961:MMT

[216] J. Ernest Wilkins, Jr. Minimum mass thin fins with specified minimum thickness. Journal of the Society for Industrial and Applied Mathematics, 9(2): 194–206, June 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gorenstein:1961:CEC

[217] Daniel Gorenstein and Neal Zierler. A class of error-correcting codes in p^{m} symbols. Journal of the Society for Industrial and Applied Mathematics, 9(2): 207–214, June 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wax:1961:PMT

[218] Nelson Wax. On a phase method for treating Sturm–Liouville equations and problems. Journal of the Society for Industrial and Applied Mathematics, 9(2): 215–232, June 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ponstein:1961:MDN

[219] J. Ponstein. Matrix description of networks. Journal of the Society for Industrial and Applied Mathematics, 9(2): 233–268, June 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Richards:1961:RVL

[220] Paul B. Richards. Representation by vortex-like singularities of obstacle flow in a straight channel. Journal of the Society for Industrial and Applied Mathematics, 9(2):269–281, June 1961. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Salzer:1961:IFN

[221] Herbert E. Salzer. An interpolation formula for “Nearly-Odd” functions, with an application to the summation of even functions. Journal of the Society for Industrial and Applied Mathematics, 9(2): 282–287, June 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rall:1961:NMC

[222] L. B. Rall. Newton’s method for the characteristic value problem Ax = λBx. Journal of the Society for Industrial and Applied Mathematics, 9(2): 288–293, June 1961. CODEN JSIMAV.

ISSN 0368-4245 (print), 2168-3484 (electronic). See errata [270].

OCallaghan:1961:DCP

[223] T. O’Callaghan. The design of a computer program to determine the natural frequencies and normal modes of vibration of an in-line mechanical system of springs and masses. Journal of the Society for Industrial and Applied Mathematics, 9(2):294–310, June 1961. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wouk:1961:OUF

[224] Arthur Wouk. On optimum utilization of faulty channels. Journal of the Society for Industrial and Applied Mathematics, 9(2):311–316, June 1961. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Greville:1961:ENF

[225] T. N. E. Greville. Erratum: “Note on Fitting Functions of Several Independent Variables” [J. SIAM 9 (1961), no. 1, 109–115]. Journal of the Society for Industrial and Applied Mathematics, 9 (2):317, June 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [208].

Butler:1961:PSS

[226] John B. Butler, Jr. On the perturbation series solution for the deflection of a semi-infinite vibrating bar on an elastic foundation. Journal of the Society for Industrial and Applied Mathematics, 9(3):319–333, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Parr:1961:ULB

[227] Wallace E. Parr. Upper and lower bounds for the capacitance of the regular solids. Journal of the Society for Industrial and Applied Mathematics, 9 (3):334–386, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Foster:1961:AWK

[228] Manus Foster. An application of the Wiener-Kolmogorov smoothing theory to matrix inversion. Journal of the Society for Industrial and Applied Mathematics, 9(3):387–392, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Moore:1961:NAE

[229] John T. Moore. Note on asymptotic expansions of modified Bessel functions. Journal of the Society for Industrial and Applied Mathematics, 9(3):393–394, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Corrsin:1961:CLS

[230] S. Corrsin and O. M. Phillips. Contour length and surface area of multiplevalued random variables. Journal of the Society for Industrial and Applied Mathematics, 9(3):395–404, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Trench:1961:GCD

[231] William F. Trench. A general class of discrete time-invariant filters. Journal of the Society for Industrial and Applied Mathematics, 9(3):405–421, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Melzak:1961:PMC

[232] Z. A. Melzak. Plane motion with curvature limitations. Journal of the Society for Industrial and Applied Mathematics, 9(3):422–432, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Douglas:1961:III

[233] Jim Douglas, Jr. On incomplete iteration for implicit parabolic difference equations. Journal of the Society for Industrial and Applied Mathematics, 9 (3):433–439, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Berman:1961:GNC

[234] Gerald Berman. A generalization of Nyquist’s criterion to the stability of active linear systems. Journal of the Society for Industrial and Applied Mathematics, 9(3):440–454, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Tu:1961:LVB

[235] Yih-O Tu and George Handelman. Lateral vibrations of a beam under initial linear axial stress. Journal of the Society for Industrial and Applied Mathematics, 9(3):455–473, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

VanGelder:1961:MTM

[236] A. Van Gelder, E. Beltram, and H. Munick. On minimum-time minimum-fuel rendezvous. Journal of the Society for Industrial and Applied Mathematics, 9 (3):474–480, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wolfe:1961:ACP

[237] Philip Wolfe. Accelerating the cutting plane method for nonlinear programming. Journal of the Society for Industrial and Applied Mathematics, 9(3): 481–488, September 1961. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Beutler:1961:EAW

[238] Frederick J. Beutler. Errata: “Approximations to Wiener Optimum Filters and Predictors” [J. SIAM 8 (1960), no. 4, 662–679]. Journal of the Society for Industrial and Applied Mathematics, 9(3): 488, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [194].

Weiss:1961:ETF

[239] Lionel Weiss. Erratum: “A Test of Fit Based on the Largest Sample Spacing” [J. SIAM 8 (1960), no. 2, 295–299]. Journal of the Society for Industrial and Applied Mathematics, 9(3):488, September 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [162].

Galletly:1961:NMB

[240] G. D. Galletly, W. T. Kyner, and C. E. Moller. Numerical methods and the bending of ellipsoidal shells. Journal of the Society for Industrial and Applied Mathematics, 9(4):489–513, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rosen:1961:GPM

[241] J. B. Rosen. The gradient projection method for nonlinear programming. part II. nonlinear constraints. Journal of the Society for Industrial and Applied Mathematics, 9(4):514–532, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gilbert:1961:RPN

[242] E. N. Gilbert. Random plane networks. Journal of the Society for Industrial and Applied Mathematics, 9(4): 533–543, December 1961. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Aumann:1961:ASC

[243] Robert J. Aumann. Almost strictly competitive games. Journal of the Society for Industrial and Applied Mathematics, 9(4):544–550, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gomory:1961:MTN

[244] R. E. Gomory and T. C. Hu. Multiterminal network flows. Journal of the Society for Industrial and Applied Mathematics, 9(4):551–570, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rice:1961:ACA

[245] John R. Rice. Algorithms for Chebyshev approximation by ab^{x} + c. Journal of the Society for Industrial and Applied Mathematics, 9(4):571–583, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Heaslet:1961:DFS

[246] Max A. Heaslet and Alberta Alksne.

Diffusion from a fixed surface with a concentration-dependent coefficient. Journal of the Society for Industrial and Applied Mathematics, 9(4):584–596, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wouk:1961:DDR

[247] Arthur Wouk. On digit distributions for random variables. Journal of the Society for Industrial and Applied Mathematics, 9(4):597–603, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Pyne:1961:EPI

[248] I. B. Pyne and E. J. McCluskey, Jr. An essay on prime implicant tables. Journal of the Society for Industrial and Applied Mathematics, 9(4):604–631, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Pandit:1961:NMC

[249] S. N. N. Pandit. A new matrix calculus. Journal of the Society for Industrial and Applied Mathematics, 9(4): 632–639, December 1961. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Taussky:1961:GTL

[250] Olga Taussky. A generalization of a theorem of Lyapunov. Journal of the Society for Industrial and Applied Mathematics, 9(4):640–643, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Krall:1961:EPR

[251] Allan M. Krall. An extension and proof of the root-locus method. Journal of the Society for Industrial and Applied Mathematics, 9(4):644–653, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Mattson:1961:NTB

[252] H. F. Mattson and G. Solomon. A new treatment of bose-chaudhuri codes. Journal of the Society for Industrial and Applied Mathematics, 9(4):654–669, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rivlin:1961:UAC

[253] T. J. Rivlin and H. S. Shapiro. A unified approach to certain problems of approximation and minimization. Journal of the Society for Industrial and Applied Mathematics, 9(4):670–699, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bureau:1962:ABS

[254] F. J. Bureau and E. J. Pellicciaro.

Asymptotic behavior of the spectral matrix of the operator of elasticity. Journal of the Society for Industrial and Applied Mathematics, 10(1):1–18, March 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Liu:1962:VPA

[255] Chen-Ya Liu. A variational problem with applications to cooling fins. Journal of the Society for Industrial and Applied Mathematics, 10(1):19–29, March 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Maehly:1962:NSC

[256] Hans J. Maehly. Numerical solution of a certain transcendental equation involving exponentials (A remark on a paper of J. R. Rice). Journal of the Society for Industrial and Applied Mathematics, 10(1):30–34, March 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Williams:1962:TTP

[257] A. C. Williams. A treatment of transportation problems by decomposition. Journal of the Society for Industrial and Applied Mathematics, 10(1): 35–48, March 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Mitchell:1962:RDO

[258] Thomas P. Mitchell and Daniel L. Pope. On the relativistic damped oscillator. Journal of the Society for Industrial and Applied Mathematics, 10(1): 49–61, March 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wilkins:1962:MMT

[259] J. Ernest Wilkins, Jr. Minimum-mass thin fins and constant temperature gradients. Journal of the Society for Industrial and Applied Mathematics, 10(1): 62–73, March 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Eberlein:1962:JLM

[260] P. J. Eberlein. A Jacobi-like method for the automatic computation of eigenvalues and eigenvectors of an arbitrary matrix. Journal of the Society for Industrial and Applied Mathematics, 10(1): 74–88, March 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See errata [282].

Freiberger:1962:RAV

[261] W. Freiberger, M. Rosenblatt, and J. Van Ness. Regression analysis of vector-valued random processes. Journal of the Society for Industrial and Applied Mathematics, 10(1):89–102, March 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Crowell:1962:GLT

[262] R. H. Crowell. Graphs of linear transformations over finite fields. Journal of the Society for Industrial and Applied Mathematics, 10(1):103–112, March 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Max:1962:SSU

[263] Joel Max. Signal sets with uniform correlation properties. Journal of the Society for Industrial and Applied Mathematics, 10(1):113–118, March 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Duffin:1962:DPM

[264] R. J. Duffin. Dual programs and minimum cost. Journal of the Society for Industrial and Applied Mathematics, 10 (1):119–123, March 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Chow:1962:NSC

[265] Tse-Sun Chow and Harold Willis Milnes. Numerical solution of a class of hyperbolic-parabolic partial differential equations by boundary contraction. Journal of the Society for Industrial and Applied Mathematics, 10(1): 124–148, March 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rice:1962:CAE

[266] John R. Rice. Chebyshev approximation by exponentials. Journal of the Society for Industrial and Applied Mathematics, 10(1):149–161, March 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wilkinson:1962:EAE

[267] J. H. Wilkinson. Error analysis of eigenvalue techniques based on orthogonal transformations. Journal of the Society for Industrial and Applied Mathematics, 10(1):162–195, March 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Held:1962:DPA

[268] Michael Held and Richard M. Karp. A dynamic programming approach to sequencing problems. Journal of the Society for Industrial and Applied Mathematics, 10(1):196–210, March 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Marshall:1962:DCG

[269] Clifford W. Marshall. Double correspondence graphs. Journal of the Society for Industrial and Applied Mathematics, 10 (1):211–227, March 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rall:1962:ENM

[270] L. B. Rall. Errata: “Newton’s Method for the Characteristic Value Problem Ax = λBx” [J. SIAM 9 (1961), no. 2, 288-293]. Journal of the Society for Industrial and Applied Mathematics, 10 (1):228, March 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [222].

Thomee:1962:SDS

[271] Vidar Thomee. A stable difference scheme for the mixed boundary problem for a hyperbolic, first-order system in two dimensions. Journal of the Society for Industrial and Applied Mathematics, 10(2):229–245, June 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Keller:1962:ELS

[272] Herbert B. Keller and Joseph B. Keller.

Exponential-like solutions of systems of linear ordinary differential equations. Journal of the Society for Industrial and Applied Mathematics, 10(2):246– 259, June 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gomory:1962:AGL

[273] R. E. Gomory and T. C. Hu. An application of generalized linear programming to network flows. Journal of the Society for Industrial and Applied Mathematics, 10(2):260–283, June 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Baer:1962:NS

[274] Robert M. Baer and Paul Brock. Natural sorting. Journal of the Society for Industrial and Applied Mathematics, 10 (2):284–304, June 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gilmore:1962:OSA

[275] P. C. Gilmore. Optimal and suboptimal algorithms for the quadratic assignment problem. Journal of the Society for Industrial and Applied Mathematics, 10 (2):305–313, June 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Struble:1962:GPS

[276] Raimond A. Struble and John E. Fletcher. General perturbational solution of the Mathieu equation. Journal of the Society for Industrial and Applied Mathematics, 10(2):314–328, June 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Osborne:1962:CBM

[277] Elmer E. Osborne. Computations of bending modes and mode shapes of single and double beams. Journal of the Society for Industrial and Applied Mathematics, 10(2):329–338, June 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wachspress:1962:OAD

[278] E. L. Wachspress. Optimum alternatingdirection-implicit iteration parameters for a model problem. Journal of the Society for Industrial and Applied Mathematics, 10(2):339–350, June 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hull:1962:CFM

[279] T. E. Hull and A. C. R. Newbery. Corrector formulas for multi-step integration methods. Journal of the Society for Industrial and Applied Mathematics, 10 (2):351–369, June 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bramble:1962:BDD

[280] J. H. Bramble and L. E. Payne. Bounds for derivatives in the Dirichlet problem for Poisson’s equation. Journal of the Society for Industrial and Applied Mathematics, 10(2):370–380, June 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Doyle:1962:DCN

[281] Worthie Doyle, J. A. McFadden, and Imanuel Marx. The distribution of a certain nonlinear functional of an Ornstein– Uhlenbeck process. Journal of the Society for Industrial and Applied Mathematics, 10(2):381–393, June 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Eberlein:1962:EJL

[282] P. J. Eberlein. Errata: “A Jacobi-like Method for the Automatic Computation of Eigenvalues and Eigenvectors” [J. SIAM 9 (1961), no. 2, 288-293]. Journal of the Society for Industrial and Applied Mathematics, 10(2):393, June 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [260].

Bather:1962:ORP

[283] J. A. Bather. Optimal regulation policies for finite dams. Journal of the Society for Industrial and Applied Mathematics, 10(3):395–423, September 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Greenfield:1962:USI

[284] Harvey Greenfield. Use of the Stodola iteration process in a Sturm–Liouville type problem with variable coefficients. Journal of the Society for Industrial and Applied Mathematics, 10(3):424– 441, September 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Levinson:1962:EPR

[285] Norman Levinson. Emission probability in a random walk. Journal of the Society for Industrial and Applied Mathematics, 10(3):442–447, September 1962. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rudd:1962:ORA

[286] Dale F. Rudd. On optimal resource allocation in multistage feedback processes. Journal of the Society for Industrial and Applied Mathematics, 10(3): 448–453, September 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Putzer:1962:RGS

[287] Eugene J. Putzer. The rate of growth of solutions of a second-order differential equation. Journal of the Society for Industrial and Applied Mathematics, 10 (3):454–468, September 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fu:1962:NRC

[288] Y. Fu and S. S. Yau. A note on the reliability of communication networks. Journal of the Society for Industrial and Applied Mathematics, 10(3): 469–474, September 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Brown:1962:NSB

[289] Robert R. Brown. Numerical solution of boundary value problems using nonuniform grids. Journal of the Society for Industrial and Applied Mathematics, 10 (3):475–495, September 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hakimi:1962:RSI

[290] S. L. Hakimi. On realizability of a set of integers as degrees of the vertices of a linear graph. I. Journal of the Society for Industrial and Applied Mathematics, 10 (3):496–506, September 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wong:1962:PES

[291] E. Wong and J. B. Thomas. On polynomial expansions of second-order distributions. Journal of the Society for Industrial and Applied Mathematics, 10 (3):507–516, September 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Miracle:1962:AST

[292] Chester L. Miracle. Approximate solutions of the telegrapher’s equation by difference-equation methods. Journal of the Society for Industrial and Applied Mathematics, 10(3):517–527, September 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Blattner:1962:BM

[293] J. W. Blattner. Bordered matrices.

Journal of the Society for Industrial and Applied Mathematics, 10(3):528– 536, September 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rosenblatt:1962:OMC

[294] M. Rosenblatt and D. Slepian. N th order Markov chains with every N variables independent. Journal of the Society for Industrial and Applied Mathematics, 10(3):537–549, September 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bellman:1962:MPM

[295] Richard Bellman and William Karush. Mathematical programming and the maximum transform. Journal of the Society for Industrial and Applied Mathematics, 10(3):550–567, September 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rosenstock:1962:ERW

[296] Herbert B. Rosenstock. Errata: “Random Walks with Spontaneous Emission” [J. SIAM 9 (1961), no. 2, 169–188]. Journal of the Society for Industrial and Applied Mathematics, 10(3):567, September 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [214].

deSanto:1962:NST

[297] Daniel F. de Santo and Herbert B. Keller. Numerical studies of transition from laminar to turbulent flow over a flat plate. Journal of the Society for Industrial and Applied Mathematics, 10 (4):569–595, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Reiss:1962:SBT

[298] Edward L. Reiss. Symmetric bending of thick circular plates. Journal of the Society for Industrial and Applied Mathematics, 10(4):596–609, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lees:1962:ADM

[299] Milton Lees. Alternating direction methods for hyperbolic differential equations. Journal of the Society for Industrial and Applied Mathematics, 10(4):

610–616, December 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Atta:1962:NEN

[300] Susie E. Atta and John A. Harvey. Numerical estimation of neutron resonance parameters. Journal of the Society for Industrial and Applied Mathematics, 10 (4):617–631, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cord:1962:PCS

[301] Marian S. Cord and Richard J.

Sylvester. The property of crosssymmetry. Journal of the Society for Industrial and Applied Mathematics, 10 (4):632–637, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Jones:1962:SSI

[302] E. E. Jones. Some solutions of an internal-external problem of the third type for a circle. Journal of the Society for Industrial and Applied Mathematics, 10(4):638–656, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

McAuley:1962:MRC

[303] Van A. McAuley. A method for the real and complex roots of a polynomial. Journal of the Society for Industrial and Applied Mathematics, 10(4): 657–667, December 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Reid:1962:FVC

[304] Walter P. Reid. Free vibrations of a circular plate. Journal of the Society for Industrial and Applied Mathematics, 10 (4):668–674, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Petryshyn:1962:GOM

[305] W. V. Petryshyn. The generalized overrelaxation method for the approximate solution of operator equations in Hilbert space. Journal of the Society for Industrial and Applied Mathematics, 10(4): 675–690, December 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Takacs:1962:CMT

[306] Lajos Takacs. A combinatorial method in the theory of queues. Journal of the Society for Industrial and Applied Mathematics, 10(4):691–694, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Tang:1962:RPB

[307] C. L. Tang. On the radiation pattern of a base-driven antenna over a circular conducting screen. Journal of the Society for Industrial and Applied Mathematics, 10(4):695–708, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See errata [397].

Curtiss:1962:IHP

[308] J. H. Curtiss. Interpolation by harmonic polynomials. Journal of the Society for Industrial and Applied Mathematics, 10 (4):709–736, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Chern:1962:GQD

[309] Shiing-Shen Chern. Geometry of a quadratic differential form. Journal of the Society for Industrial and Applied Mathematics, 10(4):751–755, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

McVittie:1962:CCT

[310] G. C. McVittie. The cosmical constant and the time-scale. Journal of the Society for Industrial and Applied Mathematics, 10(4):756–767, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Mathews:1962:GMR

[311] Jon Mathews. Gravitational multipole radiation. Journal of the Society for Industrial and Applied Mathematics, 10 (4):768–780, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sandage:1962:OTW

[312] Allan Sandage. Observational tests of world models. Journal of the Society for Industrial and Applied Mathematics, 10 (4):781–794, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Schiff:1962:GRT

[313] L. I. Schiff. General relativity: Theory and experiment. Journal of the Society for Industrial and Applied Mathematics, 10(4):795–801, December 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Konheim:1963:GCT

[314] Alan G. Konheim. A geometric convergence theorem for the perceptron. Journal of the Society for Industrial and Applied Mathematics, 11(1):1–14, March 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rice:1963:ACC

[315] J. R. Rice. Approximation with convex constraints. Journal of the Society for Industrial and Applied Mathematics, 11 (1):15–32, March 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bather:1963:ORD

[316] J. A. Bather. The optimal regulation of dams in continuous time. Journal of the Society for Industrial and Applied Mathematics, 11(1):33–63, March 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Barlow:1963:ORW

[317] Richard E. Barlow, Larry C. Hunter, and Frank Proschan. Optimum redundancy when components are subject to two kinds of failure. Journal of the Society for Industrial and Applied Mathematics, 11(1):64–73, March 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Stone:1963:MBE

[318] Jeremy J. Stone. Multiple-burst error correction with the Chinese Remainder Theorem. Journal of the Society for Industrial and Applied Mathematics, 11 (1):74–81, March 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Williams:1963:MVL

[319] A. C. Williams. Marginal values in linear programming. Journal of the Society for Industrial and Applied Mathematics, 11(1):82–94, March 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ahlberg:1963:CPS

[320] J. H. Ahlberg and E. N. Nilson. Convergence properties of the spline fit. Journal of the Society for Industrial and Applied Mathematics, 11(1):95–104, March 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Erdelyi:1963:SPB

[321] A. Erdélyi. Singular perturbations of boundary value problems involving ordinary differential equations. Journal of the Society for Industrial and Applied Mathematics, 11(1):105–116, March 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Berman:1963:DSS

[322] Gerald Berman and C. H. Cook. Determination of the stable states of active linear systems with n parameters. Journal of the Society for Industrial and Applied Mathematics, 11(1):117– 134, March 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hakimi:1963:RSI

[323] S. L. Hakimi. On realizability of a set of integers as degrees of the vertices of a linear graph II. uniqueness. Journal of the Society for Industrial and Applied Mathematics, 11(1):135– 147, March 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See note [467].

Ruiz:1963:UCC

[324] Albert L. Ruiz. Unstable characteristics of a class of functions defined by a linear differential equation with periodic coefficients. Journal of the Society for Industrial and Applied Mathematics, 11 (1):148–158, March 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

deBoor:1963:CAA

[325] Carl de Boor and John R. Rice. Chebyshev approximation by a ∏ \frac{x−r_{i}}{x+s_{i}} and application to Adi iteration. Journal of the Society for Industrial and Applied Mathematics, 11(1):159–169, March 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Guderley:1963:ESE

[326] K. G. Guderley and Mary D. Lum. On the evaluation of strongly enlarged photographs. Journal of the Society for Industrial and Applied Mathematics, 11 (1):170–182, March 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Dulmage:1963:TAB

[327] A. L. Dulmage and N. S. Mendelsohn. Two algorithms for bipartite graphs. Journal of the Society for Industrial and Applied Mathematics, 11(1):183– 194, March 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Douglas:1963:PCM

[328] Jim Douglas, Jr. and B. F. Jones, Jr. On predictor-corrector methods for nonlinear parabolic differential equations. Journal of the Society for Industrial and Applied Mathematics, 11(1):195– 204, March 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wolfe:1963:TRD

[329] Philip Wolfe. A technique for resolving degeneracy in linear programming. Journal of the Society for Industrial and Applied Mathematics, 11(2): 205–211, June 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Samson:1963:TBFa

[330] E. W. Samson and L. Calabi. On the theory of Boolean formulas: Minimal including sums, I. Journal of the Society for Industrial and Applied Mathematics, 11(2):212–223, June 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Samson:1963:TBFb

[331] E. W. Samson and L. Calabi. On the theory of Boolean formulas: Minimal including sums, II. Journal of the Society for Industrial and Applied Mathematics, 11(2):224–234, June 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kelly:1963:SRM

[332] E. J. Kelly, D. H. Lyons, and W. L.

Root. The sensitivity of radiometric measurements. Journal of the Society for Industrial and Applied Mathematics, 11(2):235–257, June 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Harrison:1963:DSM

[333] B. Kent Harrison. A discussion of some mathematical techniques used in Kron’s method of tearing. Journal of the Society for Industrial and Applied Mathematics, 11(2):258–280, June 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Henrici:1963:BEC

[334] Peter Henrici. Bounds for eigenvalues of certain tridiagonal matrices. Journal of the Society for Industrial and Applied Mathematics, 11(2):281–290, June 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See errata [440].

Karp:1963:FDS

[335] Richard M. Karp. Functional decomposition and switching circuit design. Journal of the Society for Industrial and Applied Mathematics, 11(2):291–335, June 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Imhof:1963:RQF

[336] J. P. Imhof. Remarks on quadrature formulas. Journal of the Society for Industrial and Applied Mathematics, 11(2): 336–341, June 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Aissen:1963:QDA

[337] Michael I. Aissen. Quasi-dimensional analysis and approximation formulae. Journal of the Society for Industrial and Applied Mathematics, 11(2):342– 354, June 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Srinivasan:1963:KME

[338] S. K. Srinivasan. On Katz’ model for emulsion polymerization. Journal of the Society for Industrial and Applied Mathematics, 11(2):355–359, June 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gersch:1963:MSC

[339] Will Gersch and Robert Ash. Multiple signal counting. Journal of the Society for Industrial and Applied Mathematics, 11(2):360–397, June 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cameron:1963:ANM

[340] I. G. Cameron. The application of the Von Neumann method to a steady onedimensional detonation wave. Journal of the Society for Industrial and Applied Mathematics, 11(2):398–412, June 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Stern:1963:UNT

[341] M. Stern and M. Stippes. A unified numerical treatment for vibrating beams. Journal of the Society for Industrial and Applied Mathematics, 11(2): 413–420, June 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Roos:1963:ASE

[342] B. W. Roos and W. C. Sangren. Asymptotic solutions and an equiconvergence theorem for a pair of first-order differential equations. Journal of the Society for Industrial and Applied Mathematics, 11 (2):421–430, June 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Marquardt:1963:ALS

[343] Donald W. Marquardt. An algorithm for least-squares estimation of nonlinear parameters. Journal of the Society for Industrial and Applied Mathematics, 11 (2):431–441, June 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Desoer:1963:NP

[344] C. A. Desoer and B. H. Whalen. A note on pseudoinverses. Journal of the Society for Industrial and Applied Mathematics, 11(2):442–447, June 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hansen:1963:CJM

[345] Eldon R. Hansen. On cyclic Jacobi methods. Journal of the Society for Industrial and Applied Mathematics, 11 (2):448–459, June 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Brown:1963:SRJ

[346] William M. Brown. Sampling with random jitter. Journal of the Society for Industrial and Applied Mathematics, 11 (2):460–473, June 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Buchanan:1963:NSCa

[347] Mary Louise Buchanan. A necessary and sufficient condition for stability of difference schemes for second-order initial value problems. Journal of the Society for Industrial and Applied Mathematics, 11(2):474–501, June 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Neuts:1963:MSG

[348] Marcel F. Neuts. A multistage search game. Journal of the Society for Industrial and Applied Mathematics, 11(2):

502–507, June 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Grau:1963:GBP

[349] A. A. Grau. A generalization of the Bairstow process. Journal of the Society for Industrial and Applied Mathematics, 11(2):508–519, June 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Calabi:1963:TBF

[350] L. Calabi. On the theory of Boolean formulas: Construction of the set of the formulas of minimal cost. Journal of the Society for Industrial and Applied Mathematics, 11(3):521–525, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

John:1963:SSQ

[351] Floyd I. John. Single server queues with dependent service and inter-arrival times. Journal of the Society for Industrial and Applied Mathematics, 11(3): 526–534, September 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sinden:1963:DCP

[352] Frank W. Sinden. Duality in convex programming and in projective space. Journal of the Society for Industrial and Applied Mathematics, 11(3):535– 552, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Haseltine:1963:ETNa

[353] W. R. Haseltine. Existence theorems for nonlinear ballistics. Journal of the Society for Industrial and Applied Mathematics, 11(3):553–563, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See errata [439].

Landau:1963:AAS

[354] H. G. Landau. The accuracy of the analog solution of heat conduction problems. Journal of the Society for Industrial and Applied Mathematics, 11(3): 564–578, September 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Warga:1963:CPC

[355] J. Warga. A convergent procedure for convex programming. Journal of the Society for Industrial and Applied Mathematics, 11(3):579–587, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Warga:1963:MCC

[356] J. Warga. Minimizing certain convex functions. Journal of the Society for Industrial and Applied Mathematics, 11 (3):588–593, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Warga:1963:CPS

[357] J. Warga. A convergent procedure for solving the thermo-chemical equilibrium problem. Journal of the Society for Industrial and Applied Mathematics, 11 (3):594–606, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gurtin:1963:RTL

[358] M. E. Gurtin and Eli Sternberg. A reciprocal theorem in the linear theory of anisotropic viscoelastic solids. Journal of the Society for Industrial and Applied Mathematics, 11(3):607–613, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Thacher:1963:NID

[359] Henry C. Thacher, Jr. Numerical integration and differentiation of functions of several independent variables by an iterative procedure. Journal of the Society for Industrial and Applied Mathematics, 11(3):614–622, September 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Nowinski:1963:MVT

[360] J. L. Nowinski. A mean value theorem for an arbitrary steady-state thermoelastic problem for a solid sphere. Journal of the Society for Industrial and Applied Mathematics, 11(3):623–631, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wilkins:1963:DIS

[361] J. Ernest Wilkins, Jr. Diffusion in an infinite or semi-infinite medium whose diffusion coefficient depends linearly on the concentration. Journal of the Society for Industrial and Applied Mathematics, 11 (3):632–644, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Duffin:1963:DRM

[362] R. J. Duffin and D. Hazony. The degree of a rational matrix function. Journal of the Society for Industrial and Applied Mathematics, 11(3):645–658, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Struble:1963:DDP

[363] Raimond A. Struble. A discussion of the Duffing problem. Journal of the Society for Industrial and Applied Mathematics, 11(3):659–666, September 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ben-Israel:1963:CTG

[364] A. Ben-Israel and A. Charnes. Contributions to the theory of generalized inverses. Journal of the Society for Industrial and Applied Mathematics, 11(3): 667–699, September 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Krall:1963:CER

[365] Allan M. Krall. A closed expression for the root locus method. Journal of the Society for Industrial and Applied Mathematics, 11(3):700–704, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lehmer:1963:CRS

[366] D. H. Lehmer. The complete rootsquaring method. Journal of the Society for Industrial and Applied Mathematics, 11(3):705–717, September 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Morrison:1963:ESP

[367] J. A. Morrison. The exact solution to a planar orbit problem. Journal of the Society for Industrial and Applied Mathematics, 11(3):718–731, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Mattson:1963:NAF

[368] H. F. Mattson. A note on the ambiguity function. Journal of the Society for Industrial and Applied Mathematics, 11 (3):732–736, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Descloux:1963:NCP

[369] Jean Descloux. Note on convex programming. Journal of the Society for Industrial and Applied Mathematics, 11 (3):737–747, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Olver:1963:EBF

[370] F. W. J. Olver. Error bounds for first approximations in turning-point problems. Journal of the Society for Industrial and Applied Mathematics, 11(3): 748–772, September 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lehman:1963:WDT

[371] Alfred Lehman. Wye-delta transformation in probabilistic networks. Journal of the Society for Industrial and Applied Mathematics, 11(3):773–805, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Harrison:1963:NTS

[372] Michael A. Harrison. The number of transitivity sets of Boolean functions. Journal of the Society for Industrial and Applied Mathematics, 11(3):806– 828, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Vinti:1963:ESC

[373] John P. Vinti and Tadeusz Leser. Errata: “The Sums of Certain Series Involving Bessel Functions” [J. SIAM 5 (1957), no. 1, 15–31]. Journal of the Society for Industrial and Applied Mathematics, 11(3):829, September 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [58].

Hammersley:1963:PFB

[374] J. M. Hammersley and R. S. Walters. Percolation and fractional branching processes. Journal of the Society for Industrial and Applied Mathematics, 11 (4):831–839, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Reiss:1963:EIP

[375] Edward L. Reiss. Extension of an infinite plate with a circular hole. Journal of the Society for Industrial and Applied Mathematics, 11(4):840–854, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Witzgall:1963:AIP

[376] Christoph Witzgall. An all-integer programming algorithm with parabolic constraints. Journal of the Society for Industrial and Applied Mathematics, 11 (4):855–871, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Jewell:1963:MET

[377] William S. Jewell. Multiple entries in traffic. Journal of the Society for Industrial and Applied Mathematics, 11(4): 872–885, December 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Higgins:1963:IIG

[378] Theodore P. Higgins. An inversion integral for a Gegenbauer transformation. Journal of the Society for Industrial and Applied Mathematics, 11(4): 886–893, December 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Frisch:1963:PPR

[379] H. L. Frisch and J. M. Hammersley.

Percolation processes and related topics. Journal of the Society for Industrial and Applied Mathematics, 11(4): 894–918, December 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Buchanan:1963:NSCb

[380] Mary Louise Buchanan. A necessary and sufficient condition for stability of difference schemes for initial value problems. Journal of the Society for Industrial and Applied Mathematics, 11(4): 919–935, December 1963. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Petrich:1963:NOS

[381] Mario Petrich. On the number of orthogonal signals which can be placed in a WT-product. Journal of the Society for Industrial and Applied Mathematics, 11(4):936–940, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Siekmann:1963:ARA

[382] J. Siekmann. Analysis of ring aerofoils of elliptic cross section: Part I: General theory. Journal of the Society for Industrial and Applied Mathematics, 11 (4):941–963, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Thomee:1963:PCT

[383] Vidar Thomee. A predictor-corrector type scheme for the mixed boundary problem for a hyperbolic first-order system in two dimensions. Journal of the Society for Industrial and Applied Mathematics, 11(4):964–975, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kellogg:1963:AAD

[384] R. B. Kellogg. Another alternatingdirection-implicit method. Journal of the Society for Industrial and Applied Mathematics, 11(4):976–979, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Weisfeld:1963:PON

[385] Morris Weisfeld. Polar orbits of near Earth satellites. Journal of the Society for Industrial and Applied Mathematics, 11(4):980–993, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wachspress:1963:EAA

[386] Eugene L. Wachspress. Extended application of alternating direction implicit iteration model problem theory. Journal of the Society for Industrial and Applied Mathematics, 11(4):994–1016, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Descloux:1963:ACA

[387] Jean Descloux. Approximations in L^{p} and Chebyshev approximations. Journal of the Society for Industrial and Applied Mathematics, 11(4):1017–1026, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Carlson:1963:NME

[388] David Carlson. A note on M-matrix equations. Journal of the Society for Industrial and Applied Mathematics, 11 (4):1027–1033, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Dulmage:1963:CEI

[389] A. L. Dulmage and N. S. Mendelsohn.

The characteristic equation of an imprimitive matrix. Journal of the Society for Industrial and Applied Mathematics, 11(4):1034–1045, December 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Griffin:1963:NSP

[390] D. S. Griffin and R. S. Varga. Numerical solution of plane elasticity problems. Journal of the Society for Industrial and Applied Mathematics, 11(4): 1046–1062, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See remark [555].

Kelly:1963:QPR

[391] R. J. Kelly and W. A. Thompson, Jr.

Quadratic programming in real Hilbert spaces. Journal of the Society for Industrial and Applied Mathematics, 11 (4):1063–1070, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Haseltine:1963:ETNb

[392] William R. Haseltine. Existence theorems for nonlinear ballistics part II: Effect of gravity. Journal of the Society for Industrial and Applied Mathematics, 11(4):1071–1077, December 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Barlow:1963:OCP

[393] Richard E. Barlow, Larry C. Hunter, and Frank Proschan. Optimum checking procedures. Journal of the Society for Industrial and Applied Mathematics, 11(4):1078–1095, December 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Clarke:1963:CBL

[394] S. Clarke, A. Krikorian, and J. Rausen. Computing the N best loopless paths in a network. Journal of the Society for Industrial and Applied Mathematics, 11 (4):1096–1102, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Dulmage:1963:RSO

[395] A. L. Dulmage and N. S. Mendelsohn. Remarks on solutions of the optimal assignment problem. Journal of the Society for Industrial and Applied Mathematics, 11(4):1103–1109, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Keller:1963:FAN

[396] Joseph B. Keller. The field of an antena near the center of a large circular disk. Journal of the Society for Industrial and Applied Mathematics, 11(4): 1110–1112, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Tang:1963:ERP

[397] C. L. Tang. Errata: “On the Radiation Pattern of a Base-Driven Antenna Over a Circular Conducting Screen” [J. SIAM 10 (1962), no. 4, 695–708]. Journal of the Society for Industrial and Applied Mathematics, 11(4):1113, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [307].

Hockney:1964:SLE

[398] R. W. Hockney. A solution of Laplace’s equation for a round hole in a square peg. Journal of the Society for Industrial and Applied Mathematics, 12(1): 1–14, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Erdelyi:1964:IEI

[399] A. Erdélyi. An integral equation involving Legendre functions. Journal of the Society for Industrial and Applied Mathematics, 12(1):15–30, March 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Heller:1964:LPE

[400] I. Heller. On linear programs equivalent to the transportation program. Journal of the Society for Industrial and Applied Mathematics, 12(1):31–42, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Jones:1964:FID

[401] G. Stephen Jones. Fundamental inequalities for discrete and discontinuous functional equations. Journal of the Society for Industrial and Applied Mathematics, 12(1):43–57, March 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hochstadt:1964:SCS

[402] Harry Hochstadt. On the stability of certain second order differential equations. Journal of the Society for Industrial and Applied Mathematics, 12(1): 58–59, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Putzer:1964:CSN

[403] Eugene J. Putzer. On certain systems of nonlinear differential equations. Journal of the Society for Industrial and Applied Mathematics, 12(1):60–73, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Shah:1964:SAM

[404] B. V. Shah, R. J. Buehler, and O. Kempthorne. Some algorithms for minimizing a function of several variables. Journal of the Society for Industrial and Applied Mathematics, 12(1): 74–92, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Pierre:1964:RFA

[405] Donald A. Pierre. A rational fraction approximation formula for exp[−(sD)^{1/2}] with applications. Journal of the Society for Industrial and Applied Mathematics, 12(1):93–104, March 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bramble:1964:SMV

[406] J. H. Bramble and L. E. Payne. Some mean value theorems in elastostatics.

Journal of the Society for Industrial and Applied Mathematics, 12(1):105– 114, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Turyn:1964:LGL

[407] Richard J. Turyn. The linear generation of the Legendre sequence. Journal of the Society for Industrial and Applied Mathematics, 12(1):115–116, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Campbell:1964:CST

[408] L. L. Campbell. A comparison of the sampling theorems of Kramer and Whittaker. Journal of the Society for Industrial and Applied Mathematics, 12 (1):117–130, March 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Harper:1964:OAN

[409] L. H. Harper. Optimal assignments of numbers to vertices. Journal of the Society for Industrial and Applied Mathematics, 12(1):131–135, March 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Heaslet:1964:TDC

[410] Max A. Heaslet and Franklyn B. Fuller.

Temperature distribution on conducting cylindrical shells including the effects of thermal radiation. Journal of the Society for Industrial and Applied Mathematics, 12(1):136–155, March 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rosen:1964:MBS

[411] J. B. Rosen. Minimum and basic solutions to singular linear systems. Journal of the Society for Industrial and Applied Mathematics, 12(1):156– 162, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Aziz:1964:BSS

[412] A. K. Aziz and B. E. Hubbard. Bounds for the solution of the Sturm–Liouville problem with application to finite difference methods. Journal of the Society for Industrial and Applied Mathematics, 12 (1):163–178, March 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Neuts:1964:IMO

[413] Marcel F. Neuts. An inventory model with an optional time lag. Journal of the Society for Industrial and Applied Mathematics, 12(1):179–185, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fleischer:1964:OR

[414] I. Fleischer and A. Kooharian. Optimization and recursion. Journal of the Society for Industrial and Applied Mathematics, 12(1):186–188, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Thompson:1964:DAI

[415] Robert J. Thompson. Difference approximations for inhomogeneous and quasilinear equations. Journal of the Society for Industrial and Applied Mathematics, 12(1):189–199, March 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Olver:1964:EBAb

[416] F. W. J. Olver. Error bounds for asymptotic expansions in turning-point problems. Journal of the Society for Industrial and Applied Mathematics, 12(1): 200–214, March 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rockafellar:1964:CAL

[417] R. T. Rockafellar. A combinatorial algorithm for linear programs in the general mixed form. Journal of the Society for Industrial and Applied Mathematics, 12 (1):215–225, March 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fan:1964:OMF

[418] Liang-Tseng Fan and Chiu-Sen Wang.

On the optimization of multistage feedback processes. Journal of the Society for Industrial and Applied Mathematics, 12(1):226–232, March 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cannon:1964:NSD

[419] J. R. Cannon. The numerical solution of the Dirichlet problem for Laplace’s equation by linear programming. Journal of the Society for Industrial and Applied Mathematics, 12 (1):233–237, March 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kaiser:1964:MDE

[420] Henry F. Kaiser. A method for determining eigenvalues. Journal of the Society for Industrial and Applied Mathematics, 12(1):238–248, March 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kolodner:1964:FOO

[421] I. I. Kolodner, J. Lehner, and G. M. Wing. On first order ordinary differential equations arising in diffusion problems. Journal of the Society for Industrial and Applied Mathematics, 12(2): 249–269, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cannon:1964:EES

[422] J. R. Cannon. Error estimates for some unstable continuation problems. Journal of the Society for Industrial and Applied Mathematics, 12(2):270–284, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Harrison:1964:CBF

[423] Michael A. Harrison. On the classification of Boolean functions by the general linear and affine groups. Journal of the Society for Industrial and Applied Mathematics, 12(2):285–299, June 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Moeller:1964:SIT

[424] J. W. Moeller. Signals impervious to time limited noise. Journal of the Society for Industrial and Applied Mathematics, 12(2):300–310, June 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Tee:1964:NTS

[425] G. J. Tee. A new technique for solving elliptic partial differential equations. Journal of the Society for Industrial and Applied Mathematics, 12(2):311– 347, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gomory:1964:SCN

[426] R. E. Gomory and T. C. Hu. Synthesis of a communication network. Journal of the Society for Industrial and Applied Mathematics, 12(2):348–369, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Inselmann:1964:HTG

[427] Edmund H. Inselmann. Hypothesis testing of Gaussian processes with composite alternatives. Journal of the Society for Industrial and Applied Mathematics, 12(2):370–385, June 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Newman:1964:NAD

[428] Malcolm Newman and Edward L. Reiss.

Nonlinear axisymmetric deformations of conical shells. Journal of the Society for Industrial and Applied Mathematics, 12 (2):386–402, June 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sih:1964:FSR

[429] G. C. Sih. Fracture strength of a rectangular beam with surface cracks. Journal of the Society for Industrial and Applied Mathematics, 12(2):403–412, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lemke:1964:EPB

[430] C. E. Lemke and J. T. Howson, Jr. Equilibrium points of bimatrix games. Journal of the Society for Industrial and Applied Mathematics, 12(2):413–423, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kizner:1964:NMF

[431] William Kizner. A numerical method for finding solutions of nonlinear equations. Journal of the Society for Industrial and Applied Mathematics, 12(2): 424–428, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Crout:1964:MCM

[432] Prescott D. Crout. A method for combining mixed data on position and acceleration in estimating velocities. Journal of the Society for Industrial and Applied Mathematics, 12(2):429–445, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hanson:1964:DSD

[433] M. A. Hanson. Duality and self-duality in mathematical programming. Journal of the Society for Industrial and Applied Mathematics, 12(2):446–449, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Efron:1964:OEV

[434] Bradley Efron. Optimum evasion versus systematic search. Journal of the Society for Industrial and Applied Mathematics, 12(2):450–457, June 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lindsey:1964:IIC

[435] William C. Lindsey. Infinite integrals containing Bessel function products. Journal of the Society for Industrial and Applied Mathematics, 12(2):

458–464, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

deSilva:1964:ETS

[436] C. N. de Silva. The effect of transverse shear deformation on the bending of elliptic toroidal shells. Journal of the Society for Industrial and Applied Mathematics, 12(2):465–476, June 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Pierce:1964:PBF

[437] William H. Pierce. Probability bounds in finite-state transition chains. Journal of the Society for Industrial and Applied Mathematics, 12(2):477–486, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sandberg:1964:SSL

[438] I. W. Sandberg. On the stability of solutions of linear differential equations with periodic coefficients. Journal of the Society for Industrial and Applied Mathematics, 12(2):487–496, June 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Haseltine:1964:EET

[439] W. R. Haseltine. Errata: “Existence Theorems for Nonlinear Ballistics” [J. SIAM 11 (1963), no. 3, 553–563]. Journal of the Society for Industrial and Applied Mathematics, 12(2):497, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [353].

Henrici:1964:EBE

[440] Peter Henrici. Errata: “Bounds for Eigenvalues of Certain Tridiagonal Matrices” [J. SIAM 11 (1963), no. 2, 281– 290]. Journal of the Society for Industrial and Applied Mathematics, 12 (2):497, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [334].

Sobczyk:1964:CNC

[441] Andrew Sobczyk. On the curtiss nonsingularity condition in harmonic polynomial interpolation. Journal of the Society for Industrial and Applied Mathematics, 12(3):499–514, September 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Trench:1964:AIF

[442] William F. Trench. An algorithm for the inversion of finite Toeplitz matrices. Journal of the Society for Industrial and Applied Mathematics, 12(3): 515–522, September 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Aggarwal:1964:Z

[443] H. R. Aggarwal. On the zeros of kz^{2}K′_{n}(kz)K′_{n}(z) − n^{2}K_{n}(kz)K_{n}(z). Journal of the Society for Industrial and Applied Mathematics, 12(3):523– 525, September 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Dolezal:1964:LEE

[444] Vaclav Dolezal. The location of eigenfrequencies of electrical networks. Journal of the Society for Industrial and Applied Mathematics, 12(3):526–538, September 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bazer:1964:MSO

[445] J. Bazer. Multiple scattering in one dimension. Journal of the Society for Industrial and Applied Mathematics, 12 (3):539–579, September 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cochran:1964:RZC

[446] James Alan Cochran. Remarks on the zeros of cross-product Bessel functions. Journal of the Society for Industrial and Applied Mathematics, 12(3): 580–587, September 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cline:1964:RGI

[447] Randall E. Cline. Representations for the generalized inverse of a partitioned matrix. Journal of the Society for Industrial and Applied Mathematics, 12 (3):588–600, September 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Higgins:1964:HFT

[448] Theodore P. Higgins. A hypergeometric function transform. Journal of the Society for Industrial and Applied Mathematics, 12(3):601–612, September 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Boyce:1964:RTV

[449] William E. Boyce and Bruce E. Goodwin. Random transverse vibrations of elastic beams. Journal of the Society for Industrial and Applied Mathematics, 12 (3):613–629, September 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kurowski:1964:GST

[450] G. J. Kurowski. On the general solution of the three-dimensional viscoelastic Navier’s equations for commutative systems. Journal of the Society for Industrial and Applied Mathematics, 12 (3):630–633, September 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Stoer:1964:CHN

[451] Josef Stoer. A characterization of Hölder norms. Journal of the Society for Industrial and Applied Mathematics, 12(3): 634–648, September 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wah:1964:CSV

[452] Thein Wah. Circular symmetric vibrations of ring-stiffened cylindrical shells. Journal of the Society for Industrial and Applied Mathematics, 12(3):649– 662, September 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Cottle:1964:NFT

[453] Richard W. Cottle. Note on a fundamental theorem in quadratic programming. Journal of the Society for Industrial and Applied Mathematics, 12(3): 663–665, September 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Erdogan:1964:SSD

[454] F. Erdogan and L. Y. Bahar. On the solution of simultaneous dual integral equations. Journal of the Society for Industrial and Applied Mathematics, 12 (3):666–675, September 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Chen:1964:IMF

[455] Wai-Kai Chen. The inversion of matrices by flow graphs. Journal of the Society for Industrial and Applied Mathematics, 12(3):676–685, September 1964. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lehman:1964:SSS

[456] Alfred Lehman. A solution of the Shannon switching game. Journal of the Society for Industrial and Applied Mathematics, 12(4):687–725, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sawhney:1964:CER

[457] Mathura D. Sawhney. The computation of eigenvalues. I. real symmetric matrices. Journal of the Society for Industrial and Applied Mathematics, 12(4): 726–733, December 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Tainiter:1964:SCP

[458] Melvin Tainiter. On stochastic counting processes associated with a semi-Markov process. Journal of the Society for Industrial and Applied Mathematics, 12(4):734–744, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wachman:1964:TPB

[459] Murray Wachman. On the two point boundary value problem, with application to celestial mechanics. Journal of the Society for Industrial and Applied Mathematics, 12(4):745–771, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

McAllister:1964:SNE

[460] G. T. McAllister. Some nonlinear elliptic partial differential equations and difference equations. Journal of the Society for Industrial and Applied Mathematics, 12(4):772–777, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Mangasarian:1964:EPB

[461] O. L. Mangasarian. Equilibrium points of bimatrix games. Journal of the Society for Industrial and Applied Mathematics, 12(4):778–780, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Butler:1964:LMG

[462] Terence Butler. Lagrange multipliers and the geometry of Hilbert space. Journal of the Society for Industrial and Applied Mathematics, 12(4):781– 793, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hanan:1964:CCM

[463] M. Hanan and F. P. Palermo. On cyclic codes for multi-phase data transmission systems. Journal of the Society for Industrial and Applied Mathematics, 12 (4):794–804, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gurtin:1964:NFC

[464] M. E. Gurtin. A note on functionals continuous with respect to an operator. Journal of the Society for Industrial and Applied Mathematics, 12(4):

805–806, December 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ehrlich:1964:BSS

[465] Louis W. Ehrlich. The block symmetric successive overrelaxation method. Journal of the Society for Industrial and Applied Mathematics, 12(4):807– 826, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rau:1964:RTS

[466] John G. Rau. Redundancy and trichotomous systems. Journal of the Society for Industrial and Applied Mathematics, 12 (4):827–837, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bottger:1964:NPH

[467] G. Bottger and H. Harders. Note on a problem by S. L. Hakimi concerning planar graphs without parallel elements [J. SIAM 11 (1963), no. 1, 135– 147]. Journal of the Society for Industrial and Applied Mathematics, 12(4): 838–839, December 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See [323].

Zahn:1964:ASR

[468] C. T. Zahn, Jr. Approximation symmetric relations by equivalence relations. Journal of the Society for Industrial and Applied Mathematics, 12(4): 840–847, December 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kellogg:1964:ADM

[469] R. B. Kellogg. An alternating direction method for operator equations. Journal of the Society for Industrial and Applied Mathematics, 12(4):848–854, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Pinkham:1964:RRL

[470] R. S. Pinkham. Random root location. Journal of the Society for Industrial and Applied Mathematics, 12(4): 855–864, December 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Duris:1964:RLM

[471] C. S. Duris. A Riemann-like method for solving hyperbolic difference equations. Journal of the Society for Industrial and Applied Mathematics, 12(4): 865–872, December 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Pinney:1964:SVM

[472] Edmund Pinney. The Schwinger variation method. Journal of the Society for Industrial and Applied Mathematics, 12 (4):873–885, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Traub:1964:LHI

[473] J. F. Traub. On Lagrange-Hermite interpolation. Journal of the Society for Industrial and Applied Mathematics, 12 (4):886–891, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

deBoor:1964:TPC

[474] Carl de Boor and John R. Rice. Tensor products and commutative matrices. Journal of the Society for Industrial and Applied Mathematics, 12(4): 892–896, December 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Roos:1964:SRF

[475] B. W. Roos. Successive recurrence functions and the partition of finite measure spaces. Journal of the Society for Industrial and Applied Mathematics, 12 (4):897–898, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Conner:1964:EPA

[476] Howard E. Conner. Extinction probabilities for age- and position-dependent branching processes. Journal of the Society for Industrial and Applied Mathematics, 12(4):899–909, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fiedler:1965:SEP

[477] Miroslav Fiedler. Some estimates of the proper values of matrices. Journal of the Society for Industrial and Applied Mathematics, 13(1):1–5, March 1965. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Heinbockel:1965:EPS

[478] J. H. Heinbockel and R. A. Struble. The existence of periodic solutions of nonlinear oscillators. Journal of the Society for Industrial and Applied Mathematics, 13(1):6–36, March 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Brown:1965:OPS

[479] Lawrence Brown. Optimal policies for a sequential decision process. Journal of the Society for Industrial and Applied Mathematics, 13(1):37–46, March 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sack:1965:ILE

[480] R. A. Sack. Interpretation of Lagrange’s expansion and its generalization to several variables as integration formulas. Journal of the Society for Industrial and Applied Mathematics, 13 (1):47–59, March 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Wurtele:1965:PPR

[481] Zivia S. Wurtele. A problem in pattern recognition. Journal of the Society for Industrial and Applied Mathematics, 13 (1):60–67, March 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Carrier:1965:AAT

[482] George F. Carrier. Analytic approximation techniques in applied mathematics. Journal of the Society for Industrial and Applied Mathematics, 13(1): 68–95, March 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Morrison:1965:AMA

[483] J. A. Morrison. Application of the method of averaging to planar orbit problems. Journal of the Society for Industrial and Applied Mathematics, 13 (1):96–117, March 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lucas:1965:SFP

[484] William F. Lucas. Solutions for fourperson games in partition function form. Journal of the Society for Industrial and Applied Mathematics, 13(1):118– 128, March 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Shapiro:1965:DAA

[485] Jesse M. Shapiro. Domains of attraction for absolute values of random variables. Journal of the Society for Industrial and Applied Mathematics, 13 (1):129–135, March 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kyner:1965:MTO

[486] W. T. Kyner. A mathematical theory of the orbits about an oblate planet. Journal of the Society for Industrial and Applied Mathematics, 13(1):136– 171, March 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sears:1965:SRT

[487] W. R. Sears. Some recollections of Theodore Von Karman. Journal of the Society for Industrial and Applied Mathematics, 13(1):174–183, March 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Burgers:1965:TML

[488] J. M. Burgers. The transportation of magnetic lines of force by a highly conducting fluid. Journal of the Society for Industrial and Applied Mathematics, 13 (1):184–205, March 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Crocco:1965:TBL

[489] Luigi Crocco. Turbulent boundary layer and mixing coefficient. Journal of the Society for Industrial and Applied Mathematics, 13(1):206–215, March 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Erdelyi:1965:ASP

[490] A. Erdélyi. Axially symmetric potentials and fractional integration. Journal of the Society for Industrial and Applied Mathematics, 13(1):216–228, March 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ferri:1965:MPS

[491] Antonio Ferri, Gino Moretti, and Simon Slutsky. Mixing processes in supersonic combustion. Journal of the Society for Industrial and Applied Mathematics, 13 (1):229–258, March 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Grad:1965:BE

[492] Harold Grad. On Boltzmann’s Htheorem. Journal of the Society for Industrial and Applied Mathematics, 13 (1):259–277, March 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lees:1965:KTD

[493] Lester Lees. Kinetic theory description of rarefied gas flow. Journal of the Society for Industrial and Applied Mathematics, 13(1):278–311, March 1965. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Prager:1965:MPT

[494] William Prager. Mathematical programming and theory of structures. Journal of the Society for Industrial and Applied Mathematics, 13(1):312– 332, March 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Reissner:1965:RSE

[495] E. Reissner and F. Y. M. Wan. Rotating shallow elastic shells of revolution. Journal of the Society for Industrial and Applied Mathematics, 13(1): 333–352, March 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Shapiro:1965:MAL

[496] N. Z. Shapiro and L. S. Shapley. Mass action laws and the Gibbs free energy function. Journal of the Society for Industrial and Applied Mathematics, 13 (2):353–375, June 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gilbert:1965:RMT

[497] E. N. Gilbert. Random minimal trees.

Journal of the Society for Industrial and Applied Mathematics, 13(2):376– 387, June 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Mullikin:1965:NIE

[498] T. W. Mullikin. A nonlinear integrodifferential equation in radiative transfer. Journal of the Society for Industrial and Applied Mathematics, 13(2): 388–410, June 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Yoeli:1965:LOS

[499] Michael Yoeli. Lattice-ordered semigroups, graphs, and automata. Journal of the Society for Industrial and Applied Mathematics, 13(2):411–422, June 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sreedharan:1965:SPD

[500] V. P. Sreedharan. A short proof of the duality theorem of linear programming. Journal of the Society for Industrial and Applied Mathematics, 13(2): 423–424, June 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Heinbockel:1965:PSD

[501] J. H. Heinbockel and R. A. Struble. Periodic solutions for differential systems with symmetries. Journal of the Society for Industrial and Applied Mathematics, 13(2):425–440, June 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Steiglitz:1965:OBC

[502] K. Steiglitz and A. J. Bernstein. Optimal binary coding of ordered numbers. Journal of the Society for Industrial and Applied Mathematics, 13(2): 441–443, June 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Birnbaum:1965:MCB

[503] Z. W. Birnbaum and J. D. Esary. Modules of coherent binary systems. Journal of the Society for Industrial and Applied Mathematics, 13(2):444–462, June 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Zemanian:1965:CIL

[504] A. H. Zemanian. A characterization of the inverse Laplace transforms of rational positive-real matrices. Journal of the Society for Industrial and Applied Mathematics, 13(2):463–468, June 1965. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Carlitz:1965:SMS

[505] L. Carlitz. Some multiple sums and binomial identities. Journal of the Society for Industrial and Applied Mathematics, 13(2):469–486, June 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kane:1965:DEM

[506] T. R. Kane and C. F. Wang. On the derivation of equations of motion. Journal of the Society for Industrial and Applied Mathematics, 13(2):487–492, June 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Chen:1965:MFG

[507] Wai-Kai Chen. On the modifications of flow graphs. Journal of the Society for Industrial and Applied Mathematics, 13 (2):493–505, June 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Moon:1965:DCR

[508] J. W. Moon. On the distribution of crossings in random complete graphs. Journal of the Society for Industrial and Applied Mathematics, 13(2):506– 510, June 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See erratum [558].

Pomraning:1965:VPL

[509] G. C. Pomraning. A variational principle for linear systems. Journal of the Society for Industrial and Applied Mathematics, 13(2):511–519, June 1965. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kalman:1965:IRD

[510] R. E. Kalman. Irreducible realizations and the degree of a rational matrix. Journal of the Society for Industrial and Applied Mathematics, 13(2): 520–544, June 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Brannen:1965:IMO

[511] J. P. Brannen. An iterative method for optimizing expectations. Journal of the Society for Industrial and Applied Mathematics, 13(2):545–554, June 1965. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Smiley:1965:LST

[512] M. F. Smiley. Langenhop’s stabilization theorem. Journal of the Society for Industrial and Applied Mathematics, 13 (2):555–557, June 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bazer:1965:DEW

[513] J. Bazer and L. Rubenfeld. Diffraction of electromagnetic waves by a circular aperture in an infinitely conducting plane screen. Journal of the Society for Industrial and Applied Mathematics, 13 (2):558–585, June 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sylvester:1965:TPB

[514] R. J. Sylvester and F. Meyer. Two point boundary problems by quasilinearization. Journal of the Society for Industrial and Applied Mathematics, 13 (2):586–602, June 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bauer:1965:NBR

[515] Louis Bauer and Edward L. Reiss. Nonlinear buckling of rectangular plates. Journal of the Society for Industrial and Applied Mathematics, 13(3):603– 626, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Coles:1965:MRD

[516] W. J. Coles. Matrix Riccati differential equations. Journal of the Society for Industrial and Applied Mathematics, 13 (3):627–634, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Banks:1965:SIF

[517] Dallas O. Banks. Some inequalities for the fundamental frequency of a nonhomogeneous membrane. Journal of the Society for Industrial and Applied Mathematics, 13(3):635–638, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Hobart:1965:ASC

[518] R. Hobart. An analytical solution to a class of nonlinear differential-difference equations. Journal of the Society for Industrial and Applied Mathematics, 13 (3):639–643, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Tyndall:1965:DTC

[519] William F. Tyndall. A duality theorem for a class of continuous linear programming problems. Journal of the Society for Industrial and Applied Mathematics, 13(3):644–666, September 1965. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Gordon:1965:NDE

[520] Paul Gordon. Nonsymmetric difference equations. Journal of the Society for Industrial and Applied Mathematics, 13 (3):667–673, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Keller:1965:NES

[521] Herbert B. Keller and Antoinette W. Wolfe. On the nonunique equilibrium states and buckling mechanism of spherical shells. Journal of the Society for Industrial and Applied Mathematics, 13 (3):674–705, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sakrison:1965:ESC

[522] David J. Sakrison. Estimation of signals containing unknown parameters: Comparison of linear and arbitrary unbiased estimates. Journal of the Society for Industrial and Applied Mathematics, 13 (3):706–719, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Weiss:1965:ADL

[523] Lionel Weiss. On the asymptotic distribution of the largest sample spacing. Journal of the Society for Industrial and Applied Mathematics, 13(3):

720–731, September 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Kreith:1965:EOP

[524] Kurt Kreith. Elliptic operators with a purely continuous spectrum. Journal of the Society for Industrial and Applied Mathematics, 13(3):732–739, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ponzo:1965:CRO

[525] Peter J. Ponzo and Nelson Wax. On certain relaxation oscillations: Asymptotic solutions. Journal of the Society for Industrial and Applied Mathematics, 13 (3):740–766, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Holt:1965:IIR

[526] John M. Holt. Integral inequalities related to non-oscillation theorems for differential equations. Journal of the Society for Industrial and Applied Mathematics, 13(3):767–794, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ferguson:1965:BSW

[527] Thomas S. Ferguson. Betting systems which minimize the probability of ruin. Journal of the Society for Industrial and Applied Mathematics, 13(3):795– 818, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Davisson:1965:PTS

[528] L. D. Davisson. Prediction of time series from finite past. Journal of the Society for Industrial and Applied Mathematics, 13(3):819–826, September 1965. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Ijiri:1965:GII

[529] Yuji Ijiri. On the generalized inverse of an incidence matrix. Journal of the Society for Industrial and Applied Mathematics, 13(3):827–836, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Reid:1965:EEP

[530] J. K. Reid and J. E. Walsh. An elliptic eigenvalue problem for a reentrant region. Journal of the Society for Industrial and Applied Mathematics, 13(3): 837–850, September 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Schindler:1965:GSA

[531] G. M. Schindler. On the geometric shape of axial-symmetric shock waves from a point source explosion in a nonhomogeneous medium. Journal of the Society for Industrial and Applied Mathematics, 13(3):851–863, September 1965. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Reiter:1965:DO

[532] Stanley Reiter and Gordon Sherman. Discrete optimizing. Journal of the Society for Industrial and Applied Mathematics, 13(3):864–889, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Oliver:1965:POP

[533] Robert M. Oliver and Gerold Pestalozzi. On a problem of optimum priority classification. Journal of the Society for Industrial and Applied Mathematics, 13 (3):890–901, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bollman:1965:SPP

[534] Dorothy A. Bollman. Some periodicity properties of transformations on vector spaces over residue class rings. Journal of the Society for Industrial and Applied Mathematics, 13(3):902–912, September 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Sack:1965:GLE

[535] R. A. Sack. Generalization of Lagrange’s expansion for functions of several implicitly defined variables. Journal of the Society for Industrial and Applied Mathematics, 13(4):913–926, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Williams:1965:SLP

[536] A. C. Williams. On stochastic linear programming. Journal of the Society for Industrial and Applied Mathematics, 13 (4):927–940, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic). See errata [557].

Jones:1965:LSC

[537] John Jones, Jr. On the Lyapunov stability criteria. Journal of the Society for Industrial and Applied Mathematics, 13 (4):941–945, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Klee:1965:PP

[538] Victor Klee. Paths on polyhedra. I.

Journal of the Society for Industrial and Applied Mathematics, 13(4):946– 956, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Fairweather:1965:NAD

[539] G. Fairweather and A. R. Mitchell. A new alternating direction method for parabolic equations in three space variables. Journal of the Society for Industrial and Applied Mathematics, 13(4): 957–965, December 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Konheim:1965:SWT

[540] Alan G. Konheim. The stationary waiting time distribution for a single-server queue. Journal of the Society for Industrial and Applied Mathematics, 13(4): 966–976, December 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Soong:1965:PSL

[541] T. T. Soong. On perturbation solutions to linear system dynamics — I. convergence. Journal of the Society for Industrial and Applied Mathematics, 13(4): 977–987, December 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

McFadden:1965:EPP

[542] J. A. McFadden. The entropy of a point process. Journal of the Society for Industrial and Applied Mathematics, 13 (4):988–994, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Lynch:1965:TPA

[543] Robert E. Lynch, John R. Rice, and Donald H. Thomas. Tensor product analysis of alternating direction implicit methods. Journal of the Society for Industrial and Applied Mathematics, 13 (4):995–1006, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Antman:1965:DSC

[544] Stuart Antman and William H. Warner.

Dynamic stability of circular rods. Journal of the Society for Industrial and Applied Mathematics, 13(4):1007–1018, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Schumitzky:1965:MPO

[545] Alan Schumitzky. Modulation products and optimum frequency conversion for an n-frequency input. Journal of the Society for Industrial and Applied Mathematics, 13(4):1019–1032, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Rohde:1965:GIP

[546] Charles A. Rohde. Generalized inverses of partitioned matrices. Journal of the Society for Industrial and Applied Mathematics, 13(4):1033–1035, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Konig:1965:NSC

[547] Heinz Konig and Armen H. Zemanian.

Necessary and sufficient conditions for a matrix distribution to have a positivereal Laplace transform. Journal of the Society for Industrial and Applied Mathematics, 13(4):1036–1040, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Windholz:1965:ASS

[548] Walter M. Windholz. Approximate solutions for some elasticity problems. Journal of the Society for Industrial and Applied Mathematics, 13(4):1041–1049, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Forsythe:1965:SVS

[549] George E. Forsythe and Gene H. Golub. On the stationary values of a seconddegree polynomial on the unit sphere. Journal of the Society for Industrial and Applied Mathematics, 13(4):1050– 1068, December 1965. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Waltmann:1965:AT

[550] William L. Waltmann and Robert J. Lambert. T -algorithm for tridiagonalization. Journal of the Society for Industrial and Applied Mathematics, 13 (4):1069–1078, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Pollack:1965:JPM

[551] Maurice Pollack. The jeep problem: The maximum rate of delivery. Journal of the Society for Industrial and Applied Mathematics, 13(4):1079–1086, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Posner:1965:SEC

[552] Edward C. Posner. Simultaneous errorcorrection and burst-error detection using binary linear cyclic codes. Journal of the Society for Industrial and Applied Mathematics, 13(4):1087–1095, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Weiner:1965:GIT

[553] Howard J. Weiner. The gradient iteration in time series analysis. Journal of the Society for Industrial and Applied Mathematics, 13(4):1096–1101, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Trench:1965:AIF

[554] William F. Trench. An algorithm for the inversion of finite Hankel matrices. Journal of the Society for Industrial and Applied Mathematics, 13(4): 1102–1107, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Bueckner:1966:RPD

[555] Hans F. Bueckner. Remark on a paper by D. S. Griffin and R. S. Varga “Numerical Solution of Plane Elasticity Problems” [J. SIAM 11 (1963), no. 4, 1046– 1062]. SIAM Journal on Applied Mathematics, 14(2):414, March 1966. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [390].

McKenna:1967:NAE

[556] James McKenna. Note on “Asymptotic Expansions of Fourier Integrals Involving Logarithmic Singularities” [J. SIAM 4 (1956), no. 1, 38–47], by A. Erdélyi. SIAM Journal on Applied Mathematics, 15(4):810–812, July 1967. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [45].

Williams:1967:ESL

[557] A. C. Williams. Errata: “On Stochastic Linear Programming” [J. SIAM 13 (1965), no. 4, 927–940]. SIAM Journal on Applied Mathematics, 15(1):228, January 1967. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [536].

Moon:1977:EDC

[558] J. W. Moon. Erratum: “On the Distribution of Crossings in Random Complete Graphs” [J. SIAM 13 (1965), no. 2, 506–510]. SIAM Journal on Applied Mathematics, 32(3):706, May 1977. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [508].